### v2

Abstract
In this study, we address the challenge of detecting “discursive communities” on X/Twitter by focusing on the role of verified users as the main content creators in online political debates. The analysis centers on three major Italian political events in 2022 – the Presidential election, a governmental crisis, and the general elections – occurring before the introduction of paid account verification. We propose and compare two novel methodologies, MonoDC and BiDC, which exploit, respectively, the retweet network among users and a similarity network based on shared audiences, while integrating a maximum entropy null model to filter out the inherent noise in online social networks. Our results demonstrate that leveraging verified users—considered as indicators of prestige and authority—leads to significantly clear community partitions that closely reflect the actual political affiliations, outperforming standard community detection algorithms applied to the entire retweet network. Moreover, the comparison of different methodologies and user sets suggests that the status conferred by the blue verification tick plays a dominant role in shaping online discourse, with important implications for platform governance, especially in light of the recent shift to paid verification.
1 Introduction
Online communities naturally emerge from users’ tendency to follow and interact with others who share similar interests, experiences, connections, or political views [1, 2, 3]. Such communities can provide valuable support and shared knowledge [4, 5, 6]; however, they can also form “echo chambers,” where members are exposed only to a narrow range of perspectives that reinforce their existing beliefs [7, 8, 9, 10, 11, 12, 13]. To fully understand the influence of user communities on information diffusion and opinion dynamics, it is crucial to develop effective methods for detecting these communities and capturing their key structural features.
In the literature, several refined approaches have been proposed to enhance standard community detection methods by tailoring them to the context of online social networks. For example, Citraro and Rossetti [14] incorporate account attributes to strengthen the conventional definition of communities. Other approaches leverage the varied roles played by different accounts in online debates. Although Twitter/X was initially designed as a many-to-many platform, numerous studies have observed that it actually functions as a few-to-many network, where a small number of influential content creators generate the majority of posts [15, 16, 17]. In this context, it becomes natural to focus on the leading content creators for detecting communities on Twitter/X.
Identifying which accounts should be considered as leading content creators is, therefore, essential. In [4], influential users—defined as those with more followers than followees and with more mentions received than given—are identified as the primary figures in the debate. Similar approaches, based on user activity, have also been explored [15, 16, 17]. Other studies place more emphasis on account prestige than on activity, considering verified users as the leading content creators [18, 19, 20].
Verified users on X/Twitter have an interesting history: they were first introduced in June 2009 to curb the impersonation of notable persons or organizations [21, 22]. In this early stage, the platform assumed responsibility for identifying key public figures and verifying their identities with a blue checkmark, while standard users were not part of the verification process [23]. Efforts to broaden the pool of verified accounts through user applications were unsuccessful [21]. Eventually, in November 2022, verification became a feature of X’s Premium accounts, available upon payment [24]—a change that has sparked concerns regarding the influence of verified users on online debates [25, 26].
In the present study, we investigate the method proposed in [18] using three distinct datasets collected during 2022, in a period before blue ticks could be purchased via X’s Premium subscription. Our framework comprises: (i) identifying verified users as the leading content creators; (ii) constructing an informative network of verified users by filtering out random noise through a comparison between empirical data and an appropriate null model; (iii) detecting communities within this sub-network; and (iv) propagating the labels to all other users. Following the terminology introduced by [27], the resulting partition is referred to as “discursive” communities.
A key component of our approach, in addition to leveraging verified users, is the elimination of random noise from the system by comparing empirical data against a maximum entropy statistical benchmark. As numerous studies have shown, removing random noise is crucial in online social networks to extract the relevant signal regarding users’ preferences and behaviors [18, 28, 29]. In this regard, maximum entropy null models provide an ideal statistical benchmark, being maximally random apart from the constraints imposed [30].
Figure 1: MonoDC’s pipeline. Starting from the data from Twitter/X, we build a bipartite directed network between accounts and posts; a directed arrow from a user to a post represents authorship, while in the opposite direction represents a retweet. Then, the content of the bipartite network is projected into a monopartite directed network of users. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite directed networks. Then, the sub-network relative to content creators is analysed, and communities are detected using standard algorithms. Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
Compared to previous work, we distinguish between two main versions of the algorithm that differ in how step (ii) is implemented. In $\mathrm{MonoDC}$ (Monopartite-network-induced Discursive Communities), we consider the direct interactions —i.e., retweets— among content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig. 1. In $\mathrm{BiDC}$ (Bipartite-network-induced Discursive Communities), we instead construct a similarity network based on the shared audience between pairs of leading content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig. 2. Additionally, we explore various classes of leading content creators beyond verified users.
Our proposed methods were tested on three large datasets linked to the primary political events in Italy during 2022, before user verification became a feature of Premium accounts. To validate our results, we manually annotated approximately 300 verified accounts to determine their political leanings and compared different clustering algorithms to assess their ability to detect this signal.
The key finding is that both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$, when using verified users as the main content creators, significantly outperform standard algorithms. They produce clearer partitions that consist of a few meaningful communities, which accurately reflect the political landscape during the studied period. Specifically, $\mathrm{MonoDC}$ appears to capture political parties more effectively—likely because politicians predominantly retweet their party leaders—while $\mathrm{BiDC}$ better identifies political alliances due to the shared construction of narratives. Importantly, our results deteriorate markedly when the set of main content creators is altered.
Figure 2: BiDC’s pipeline. Starting from the data from Twitter/X, we build a bipartite undirected network between the accounts of the leading content creators and the one of standard users; a link is present if the standard users retweeted the given content creator at least once. Then, the information of the bipartite network is projected into a monopartite network of leading content creators. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite undirected networks. Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
In summary, our study demonstrates that verified accounts play a central role in shaping discursive communities on X/Twitter. This evidence carries important implications, especially considering that our datasets are limited to the period before the implementation of X’s Premium accounts, when blue checkmarks were granted based on the platform’s autonomous verification of key accounts. While we cannot conclusively determine whether the effectiveness of verified users as leading content creators stems solely from the prestige of the verification or from other factors, our findings indicate that definitions based exclusively on user activity yield inferior performance. If the prominence of verified users in steering the online discourse is indeed due to the prestige associated with blue ticks, then introducing verification as a paid feature could pose serious challenges to the platform’s integrity, potentially enabling malevolent actors to purchase authority to distort the political agenda.
The paper is organised as follows. In Section 2, we introduce the methodology adopted in this study. Specifically, we present maximum entropy null models and their application to the detection of discursive communities (Subsections 2.1 and 2.2). We also define the V-measure [31], the metric used to compare different partitions (Subsection 2.3). Section 3 presents our results. We begin by illustrating the limitations of standard methods in reproducing non-trivial benchmarks observed in real-world data (Subsection 3.3). We then show how $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ address these issues when verified users are identified as the primary content creators (Subsection 3.4). Additionally, we show that the performance of both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ deteriorates when different groups of leading content creators are considered (Subsection 3.5). Finally, in Section 4, we summarise our findings and discuss their potential implications.
2 Methods
2.1 Maximum-entropy null models and discursive communities
As underlined by many studies across a multitude of applications  [18, 19, 20, 27, 32, 33, 34, 35, 36, 28, 29, 13], using a proper statistical validation is critical for the analysis of online social networks (OSNs) that are particularly prone to noise. Entropy-based null models [30] represent a tailored and, at the same time, unbiased tool for detecting the statistically significant signal in complex networks. The method essentially consists in: i) defining an ensemble of graphs, i.e. the set of all possible graphs having the same number of nodes as the observed network; ii) performing a constrained maximization of the Shannon entropy of the ensemble, fixing some constitutive properties of the original system [37]; iii) maximizing the (log-)likelihood of the observed network [38], to get the numerical values of the parameters of an effective fitness model.Remarkably, this framework mimics Jaynes’ derivation of statistical ensembles in Physics starting from Information Theory [39].
Let us start from a real network $G^{*}$ (in the following, all quantities related to the observed network will be denoted by an asterisk *). Firstly, let us define a set $\mathcal{G}$ – the ensemble– including all possible graphs having the same number of nodes $N$ as in $G^{*}$ [^1] Since the number of nodes is fixed in the randomization, we drop the asterisk * from $N$.. Let us identify a vector of quantities $\vec{C}$ that are fundamental for the description of the real system. A proper benchmark for $G^{*}$ would be completely random, but for the information contained in $\vec{C}(G^{*})$: in this way it is possible to highlight all properties of $G^{*}$ that are not trivially due to $\vec{C}$. Translating this rationale in maths, we need to perform a constrained maximization of the Shannon entropy (to have a maximally random benchmark) associated to the ensemble, where the constraints are exactly $\vec{C}(G^{*})$. In this sense, the (ensemble) average $\langle\vec{C}\rangle$ will be set exactly to $\vec{C}(G^{*})$.
If $S$ is the Shannon entropy, defined as
$S=-\sum_{G\in\mathcal{G}}P(G)\ln P(G),$
then we can perform the constrained maximisation using the method of Lagrange multipliers, i.e. by maximising $S^{\prime}$ defined as
$S^{\prime}=S+\vec{\theta}\cdot\Big{(}\vec{C}(G^{*})-\langle\vec{C}\rangle\Big{)}+\alpha\Big{(}1-\sum_{G\in\mathcal{G}}P(G)\Big{)},$
where $\vec{\theta}$ is the vector of Lagrange multipliers associated with $\vec{C}$ and $\alpha$ is the Lagrangian multiplier associated to the normalization of the probability $P(G)$. The probability solving the maximization problem above has the standard functional form of exponential random graphs:
$P(G)=\frac{e^{-\vec{\theta}\cdot\vec{C}(G)}}{Z(\vec{\theta})},$ (1)
where $Z=e^{\alpha+1}=\sum_{G\in\mathcal{G}}e^{-\vec{\theta}\cdot\vec{C}(G)}$ is the partition function. In order to get the numerical value of the Lagrangian multipliers $\vec{\theta}$ we have to explicitly impose
$\langle\vec{C}\rangle=\vec{C}(G^{*}).$ (2)
It can be shown that the condition in Eq. (2) can also be obtained through the maximization of the (log-)likelihood [38].
The framework described above is particularly flexible and can be adapted to different kinds of networks. In the present paper, we are interested in its bipartite versions for bipartite undirected and bipartite directed networks, i.e. respectively, the Bipartite Configuration Model (or BiCM [40]) and the Bipartite Directed Configuration Model (or BiDCM [41, 18]). In both cases, the constraints imposed in the entropy maximisation are, respectively, the undirected and directed degree sequences. More details about BiCM and BiDCM can be found in Appendices A.1 and A.2.
2.2 $\mathrm{MonoDC}$ and $\mathrm{BiDC}$
In the following, the two methods that we are proposing are presented. They both leverage the presence of leading content creators, i.e. users mostly posting original messages and less inclined to repost the messages created by others. There are multiple ways to define leading content creators: at the present level, we prefer to remain as general as possible. Therefore, in the following those users are called $\top$-⊤ -users, while the $\top$ users are called $\bot$-⊥ -users.
2.2.1 $\mathrm{MonoDC}_{\top}$
At least in the Italian political debate, it is common to observe that less relevant political figures are particularly active in liking and reposting the content created by more central leaders [^2] https://www.ilpost.it/2024/10/31/post-giorgia-meloni-social-commenti-parlamentari-fdi/. There could be multiple reasons for such a behaviour: intercept the audience of the leaders, demonstrate loyalty to either the leaders, the party or the objective of the campaign, or simply to contribute to the campaign of the leaders by increasing the visibility of their content. Therefore, it is reasonable to expect the activity of politicians to be much more focused on the activity of their party than other users.
To capture such a phenomenon, the MonoDC studies the interactions among content creators through their activity in creating new contents or reposting the one created by others. The steps of MonoDC are represented as a cartoon in Fig. 1. We describe data from Twitter/X as a bipartite directed network between users and posts in which edges from users to posts indicate authorships, while the vice versa indicates retweets. The information contained in the directed bipartite network is then projected to describe the retweet network, i.e. a directed multilink network in which nodes are users and the number of edges from user $A$ to user $B$ is the number of retweets that $B$ made to the post written by $A$. In this representation, we expect the (bipartite directed) degrees to play a role, since they represent the activity of users in either creating new posts or retweeting existing ones (represented, respectively, as the out- and in-degree in the bipartite directed network of the user nodes) and the virality of the various posts (i.e. the out-degree of the post nodes). Therefore, we can use a maximum entropy null model for directed bipartite networks as a statistical benchmark to validate the empirical retweet network, i.e. the Bipartite Directed Configuration Model (BiDCM) mentioned in Section 2.1. Technical details about BiDCM can be found in Appendix A.2.
The result of such validation is a directed binary network, in which an edge represents that the retweet activity between the two nodes is non-trivial, i.e. it cannot be explained by just the attitude of users towards posting and reposting and the virality of the messages shared. We further restrict the validated network to only users in the set of content creators, i.e. the $\top$ user, in our jargon.
The main idea is that the null model is not informed about the orientation of the users and therefore expects the debate to be “flat”, i.e. without any particular focus of the users towards the content created by some of the accounts. In this sense, communities should emerge naturally by comparing empirical data with the null model. Such a structure can be further captured by running a community detection algorithm on the validated retweet network.
Finally, the labels obtained by the community detection are extended to $\bot$ users (i.e. the users that are not in the $\top$ set) by implementing Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network restricted to $\top$-users.
2.2.2 $\mathrm{BiDC}_{\top}$
The intuition behind $\mathrm{BiDC}_{\top}$ is to capture discursive communities through the perception of “standard” users of the similarity in the content created by the leaders of the debates [18, 43, 27, 19]. Such a phenomenon is more stressed during political debates, being more or less intense depending on the electoral system of the country [43].
Following the procedure outlined in Fig. 2, we built the bipartite network between leading content creators and “standard” users: a link is present if the standard user has retweeted at least once the corresponding content creator. We then projected the bipartite network to a monopartite co-occurrences network among content creators. As in the case of $\mathrm{MonoDC}_{\top}$, we validate the network by comparing the empirical data with a maximum entropy null model for bipartite networks, i.e the Bipartite Configuration Model (BiCM) mentioned in Section 2.1. In this case, the information about the degree of the users is capturing the number of different retweeters and the number of different content creators retweeted for $\top$-⊤ - and $\bot$-⊥ -users, respectively; more details about the implemented maximum entropy null model can be found in Appendix A.1.
BiCM, while accounting for the degree sequences of both layers, is not informed about the presence of communities and therefore expects the co-occurrence network not to have a modular structure. In this sense, any non-trivial focus of the users towards groups of content creators should be captured by comparing empirical data against the null model. Such a structure can be further refined by running a community detection algorithm on the validated projection network.
Similarly to $\mathrm{MonoDC}_{\top}$, labels are extended to $\bot$ users by using Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network.
2.3 Measuring the similarity of two partitions
To assess the similarity between two ways of partitioning the same set of users, we use of the V-measure ($\mathrm{VM}_{\beta}$), a parametric measure of cluster similarity proposed by [31].
Let $S(P)$ be the Shannon entropy of $P$, and $S(P\mid Q)$ be the conditional entropy of $P$ given $Q$. If $P$ and $Q$ are two partitions, the $\mathrm{VM}_{\beta}(P,Q)$ is defined as:
$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)I(P,Q)}{S(P)+\beta S(Q)}$
where $I(P,Q)$ is the mutual information between $P$ and $Q$.
For $\beta=1$ we have
$\mathrm{VM}_{1}(P,Q)=\frac{2I(P,Q)}{S(P)+S(Q)}=1-\frac{VI(P,Q)}{S(P)+S(Q)}$
i.e., the $\mathrm{VM}_{1}$ accounts for the mutual information between the two partitions in a way that does not depend on their entropy, being equivalent to the Variation of Information [44], normalized by the total entropy of the partitions. With respect to other information-theoretic metrics, the $\mathrm{VM}_{1}$ and the VI have the advantage of being two distances – in particular, of satisfying the triangle inequality. Further, the $\mathrm{VM}_{1}$ always takes values in $[0,1]$, making it possible to directly compare the scores obtained for different combinations of algorithms and/or datasets. By letting $\beta$ vary in $[0,+\infty)$, we consider both scenarios in which one partition consists of sub-communities (e.g., political movements/trends within parties) or super-communities (e.g., political coalitions) of the other one. More details about the V-measure can be found in Appendix B.
3 Results
3.1 Data and networks
To compare our approach for discursive community with standard community detection algorithms, we collected Twitter data regarding the three main political events debated on Italian Twitter in 2022. For each such topic of discussion, we identified a set of relevant keywords that were used to collect tweets using Twitter’s Academic API. Let us remark that data were downloaded before user’s verification became a feature of the X’s Premium subscription [^3] https://en.wikipedia.org/wiki/Twitter_verification.. A summary of the datasets considered in this paper is reported in Table 1.
Table 1: A summary of the datasets considered in the present paper
Dataset Unique users Unique tweets Annotated users Verified users Influential users H-$\geq$3 users
President1 119.018 711.497 292 1148 5804 1296
Crisis2 113.876 1.414.181 229 1004 7038 2554
Elections3 222.356 2.784.951 239 1555 12184 4382
  • 1
    Debate around the election of the President of the Republic. The political debate ended with the re-election of Sergio Mattarella on January 29, 2022.
  • 2
    Debate around the political crisis that led to the fall of the Draghi government on July 21, 2022.
  • 3
    Debate around the Italian general elections held on September 25, 2022, that led to Giorgia Meloni becoming the new Prime Minister.
3.2 The benchmark
In general, there is no reason to expect data to cluster due to some metadata [45]. Multiple effects can cause –especially in online social networks– different users to interact with the same content creators. Nevertheless, due to the strictly political nature of the debates captured by our data, we expect to observe a signal of the political orientation of the users in the cluster that can be used as a benchmark for the classification methods under study.
To construct our benchmark labeling, we first focused on just the verified accounts of Italian politicians and political parties. We choose verified accounts since their identities are certified by the platform. We manually labeled these accounts according to their political party and then grouped these labels according to the coalition each party belonged to during the general elections. This procedure yielded two annotation levels, which we refer to as “parties” and “coalitions”.
To check if this benchmark is well-posed, we studied the retweet network restricted to the annotated users. Retweets are usually considered as a sort of endorsement of the content shared [1], so we expected users with a similar political orientation to cluster in this network. The modularity of the partitions induced by the annotations, whose values are all above 0.6 (cfr. Table 2), confirms our intuition and supports the idea of measuring the quality of a community detection algorithm also based on its ability to retrieve our manual annotations.
Table 2: Retweet networks of annotated users and modularity (Q) of the partition induced by the manual annotation at the party and coalition level.
Dataset Nodes Links $Q(\text{parties})$ $Q(\text{coalitions})$
President 169 219 0.763 0.618
Crisis 145 246 0.721 0.641
Elections 154 355 0.777 0.649
To compare different algorithms not just based on the small minority of users that we manually annotated, we extended the annotations to the rest of the network using a well-known Label propagation algorithm [42] –taking the most frequent label out of 100 independent realizations. These propagated labels cannot be considered equally accurate, but they arguably represent the best possible guess of a user’s political leaning, which can be made only based on how that user is connected to verified politicians in the retweet network.
Let us notice that such an approach favors our models, in principle, since we are using the same label propagation algorithm to label standard users. In a sense, if we correctly tag $\top$ users, we will implicitly also meet the benchmark for $\bot$ users. It is indeed true, but it is an honest way to propagate the labels of verified accounts, as we cannot trust, in general, the one of other accounts that could be present in the user description.
3.3 Limits of standard methods
Let us start by considering the partitions obtained running four off-the-shelf community-detection algorithms on the entire retweet network: (i) Louvain, a heuristic for modularity maximization; (ii) Label propagation, a message passing algorithm where each node tries to align with the majority of its neighbors; (iii) Infomap, which is based on finding the optimal encoding for random walks on the graph; (iv) Inference, whose rationale is inferring the parameters of a stochastic block-model through a maximum-likelihood approach.
For Louvain, Label propagation and Infomap, we used the implementation available in igraph (igraph.org, Python interface). For Inference, we used the minimize_blockmodel_dl() function provided in graph-tool (graph-tool.skewed.de), that fits a stochastic block-model to the given graph by minimizing the model’s description length. To guide the inference process, we set it explicitly to search for a planted-partition, that is, exclusively for assortative communities. In addition, we refined the algorithm as recommended in graph-tool’s documentation, that is, by running 100 times the multiflip_mcmc_sweep() method with zero temperature and 100 iterations.
We ran each algorithm 100 times [^4] Each single run of the Louvain algorithm is already obtained taking the partition with the greatest modularity out of 100 independent executions with shuffled vertices. Similarly, each run of the Label propagation algorithm is already obtained by taking, for each node, the most frequent label out of 100 independent executions. and measured: the $\mathrm{VM}_{1}$ between each partition, restricted to the annotated users, and the annotations, both at the party and coalition level; the pairwise $\mathrm{VM}_{1}$ between partitions obtained with different algorithms, restricted to the verified users. Fig. 3 reports these $\mathrm{VM}_{1}$ values, considering only the parties annotations. We focused on party annotations, but similar results are observed with other algorithms and with coalitions.
Fig. 3 highlights the limited accuracy in recognizing the affiliation of renowned politicians –$\mathrm{VM}_{1}\in[0.4,0.6]$– of the communities assigned to verified users by different off-the-shelf algorithms.
Furthermore, as shown in Fig. 10 in Appendix C, standard algorithms show great variability of the communities assigned to verified users by different off-the-shelf algorithms –pairwise $\mathrm{VM}_{1}\in[0.3,0.7]$. Overall, these findings motivate the research for a different approach to the detection of discursive communities on Twitter/X.
Figure 3: Limits of standard methods. The $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms, restricted to the annotated users, and the parties’ annotations. Standard methods show a limited accuracy in recognising the affiliation of renowned politicians.
3.4 $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$
We now show the results of our methods, i.e. $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, when we choose verified users as the class of content creators: in this case, we will write $\top=V$. In the following, the communities on the network of $V$ users are obtained by running the Louvain algorithm on the projected network and are then propagated to the rest of the network. We tested our community detection algorithm, as the Label propagation and Infomap: their results are very similar to those obtained with Louvain ($\mathrm{VM}_{1}>0.8$) and yield similar results also for both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$. Communities obtained with the inference-based method are instead way less accurate in reconstructing the annotations.
3.4.1 Comparison with standard methods
In Fig. 4, we use the $\mathrm{VM}_{1}$ to show how accurately each method reconstructs the manual annotations at the party and coalition level, respectively in the first and second row. Each violin represents the distribution over 100 independent runs, and we compare $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with the partitions obtained running the Louvain algorithm on the entire retweet network –which performs no worse than other algorithms (cfr. Fig. 3). $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions. The fact that in all three cases the community detection algorithm used is Louvain signifies that what ultimately matters is focusing on the right network.
For completeness, the comparison between the performances of our methods and Louvain is presented in Fig. 11 in Appendix C. Even in this case, both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ outperform Louvain. Nonetheless, let us note that, when comparing the results of the propagated labels of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}$ with the annotations of verified users propagated using the same label propagation algorithm, our methods have an advantage. Nevertheless, the performances of standard methods are worse, even when limited to annotated users.
Figure 4: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods. The $\mathrm{VM}_{1}$ between partitions of the annotated users, obtained with different methods, and the annotations at party (first row) and coalition (second row) level. $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions.
It could be argued that standard algorithms, when executed on the entire retweet network, return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect. Fig. 5 shows that this is not the case: when compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e. at essentially all scales at which the political spectrum can be observed.
Figure 5: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods at different resolution levels. The $\mathrm{VM}_{\beta}$ between partitions of the annotated users, obtained with different methods, and the annotations, for different values of $\beta$ and both at the parties (top) and coalitions (bottom) level. Each point is the average over 100 runs of the method indicated by the colour and marker. Analogous results are obtained when considering the propagated annotations. When compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e. at essentially all scales at which the political spectrum can be observed. In this sense, standard algorithms, when executed on the entire retweet network, do not return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect.
In Fig. 6, the Sankey diagrams corresponding to each evaluated method provide a detailed depiction of the political composition of the extracted communities. $\mathrm{MonoDC}_{V}$ exhibits remarkable precision in delineating party structures —successfully distinguishing individual parties such as Azione and Italia Viva, which presented a unified list during the elections— and further reveals latent political currents within party boundaries. Conversely, $\mathrm{BiDC}_{V}$ proves highly effective in reconstructing the official electoral coalitions, achieving a near-exact correspondence with their institutional configurations. The Louvain algorithm, applied to the entire retweet network, yields a substantially noisier community structure that conflates partisan and coalition-level affiliations. Notably, Forza Italia emerges as an isolated community, distinct from its allied parties, while several political figures are clustered within communities dominated by their ideological adversaries.
(a) Louvain on RT network.
(b) $\mathrm{MonoDC}_{V}$.
(c) $\mathrm{BiDC}_{V}$.
Figure 6: Qualitative comparison of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with Louvain on the Elections dataset. For each of the considered methods, the Sankey diagrams show the composition of the extracted communities in terms of political parties. $\mathrm{MonoDC}_{V}$ very accurately detects parties (Azione and Italia Viva ran together as a single list at the elections), plus some political movements/trends within parties; $\mathrm{BiDC}_{V}$ almost perfectly detects the official political coalitions of those elections; Louvain, executed on the entire retweet network, provides a more noisy classification that is a mix of the two levels, with Forza Italia separated from its coalitions and many politicians contributing to communities led by some of their rivals.
Finally, since $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ capture the political leaning of users at different scales, in Fig. 7 we use the $\mathrm{VM}_{\beta}$ to check whether the clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$. This seems to be true, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
Figure 7: Different resolution of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$. The $\mathrm{VM}_{\beta}$, for variable $\beta$, between partitions of all users obtained running $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$. The clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
3.4.2 Final remarks
Both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ benefit from a pre-processing stage capable of extracting, from the original retweet data, a more informative network of just such leading content creators. The role that statistical validation plays in the two cases is, however, incomparable.
In $\mathrm{MonoDC}_{V}$, the validated retweet networks of verified accounts of politicians include 94-97% –depending on the dataset– of all edges present before the validation. The validation is performed with respect to the patterns emerged from the analysis of the entire network and the fact that almost all retweets made by politicians are statistically significant means that their activity on X is extremely focused on just the content produced by a limited niche of users. Since these users are their party’s leaders and fellow members, $\mathrm{MonoDC}_{V}$ is the approach that best captures the loyalty to the party and, hence, the difference between parties that apparently occupy a similar position in the political spectrum. The price to pay for this finer classification is its minor stability, as can be seen in Fig. 4. In this sense, $\mathrm{MonoDC}_{V}$ seems particularly promising to examine the mesoscale structure of such a subgraph to infer the main poles of the political discussion, since the phenomenon we are leveraging on is particularly strong.
Differently from the $\mathrm{MonoDC}_{V}$’s case, for $\mathrm{BiDC}_{V}$ the impact of the statistical validation is critical: the number of links in the validated network is 7-12% –depending on the dataset– of those present in the non-validated projection. When the political sides are clear, this validated signal coming from the audience returns a very stable partition (cfr. Fig. 4). In addition, $\mathrm{BiDC}_{\top}$ has the advantage that the similarity between any two content creators $u,v\in\top$ can be established even when $u$ and $v$ did not directly interact, making $\mathrm{BiDC}_{\top}$ more flexible, robust and widely usable than $\mathrm{MonoDC}_{\top}$.
For the sake of completeness, we ran the same set of community-detection algorithms considered in Section 3.3 on the non-validated network, as well as on the validated projections with different significance levels [^5] To properly target the significance level in validating the projected network, we implemented the False Discovery Rate (FDR) procedure [46], see Appendices A.1.1 and A.2.1., cfr. Fig. 12 in Appendix C. The results on the non-validated network are very poor, whereas the significance level does not seem to have a great impact on the inferred similarity network.
3.5 Other classes of content creators
In the previous subsections, we examined the behaviours of the $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ after identifying verified users as relevant content creators, i.e. the main leaders of the shared discourses. Verified users ($V$) are a very natural choice for the $\top$ layer of both $\mathrm{MonoDC}_{\top}$ –because their behavior on the platform is bound to their role outside of it– and $\mathrm{BiDC}_{\top}$ –because they are generally known to and easily categorizable by the audience.
There are other ways, however, to select users based on their actual or presumed influence on online discussions. For instance, [4] defined “influential” users those who have more followers than friends (i.e. the accounts a user follows, sometimes referred to as “followees” in the literature), and who, in the debate under study, received more mentions than those they made. Another way to measure the impact of a user in a community, commonly employed for scientists and academics, is the Hirsch index [47] –shortly, H-index– which can be adapted to X’s mode of operation by saying that a user $u$’s H-index in the debate is the largest $h$ such that at least $h$ tweets authored by $u$ received at least $h$ retweets each.
In the following, we use $I$ to denote the set of influential users and $H$ for the set of top H-index users (defined as users having H-index3absent3$\geq$ 3≥ 3, see Fig. 13 in Appendix C for a justification of the threshold). We aim to assess both what happens when $\top=I$ or $\top=H$ is used in $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, and how stable and accurate is the classification obtained for these two sets of central users when compared with the entire dataset.
(a) $\mathrm{MonoDC}_{\top}$, for different choices of $\top$, compared with annotations at the parties level.
(b) $\mathrm{BiDC}_{\top}$, for different choices of $\top$, compared with annotations at the coalitions level.
Figure 8: Comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$. The $\mathrm{VM}_{1}$ between partitions obtained with different choices of $\top$ and the annotations, for the Elections dataset. For each method, we used the level of annotation that the method more accurately reproduces, that is, parties for $\mathrm{MonoDC}_{\top}$ (a) and coalitions for $\mathrm{BiDC}_{\top}$ (b). The columns refer to the user set upon which the VM is computed. For instance, “user set = influential” means that in the second column we only consider how influential users have been labeled by different methods, temporarily ignoring the rest of the network.
In Fig. 8, the $\mathrm{VM}_{1}$ is used to compare partitions obtained with different choices of $\top$ with the (propagated) annotations ($x$-axis). Here we consider the Elections dataset, but similar results hold for the other datasets.
In Appendix C), we also compare the cases $\top=H$ and $\top=I$ with the case $\top=V$, taken as a reference.
The choice of the $\top$ set has an impact on both $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$. Especially for $\mathrm{BiDC}_{\top}$, $\top=V$ yields the best overlap with the annotations. It could be argued that it is due to $V$ being a small set that includes all manually annotated users, but $H$ outperforms $I$ despite the former includes just 20-50% (depending on the dataset) of the annotated public figures, compared with 80-90% of the latter.
This supports the claim that $V$ users are the key to understanding the sides of the public discourse.
4 Conclusions
In this paper, we focused on X/Twitter and on so-called discursive communities, i.e. the sub-groups that naturally emerge in Online Social Networks (OSNs) as users cluster around shared narratives, language, or value systems during public debates. What we observe in OSNs is a network of interactions, where edges exist for a combination of factors – e.g. the visibility, the status and the behaviour of different users on and off the platform, and the platform’s recommendation system – and where understanding what led to the creation of a community is almost as important as detecting the community itself.
In this paper, we investigate the role of verified users in shaping discursive communities in the pre-Premium accounts era. Introduced in 2009, verified accounts were originally intended to prevent impersonation of public figures or organizations [21, 22]. At first, the platform itself selected which accounts to verify, based on their relevance in public discourse. These accounts were marked by a blue checkmark next to their name. While verification helped authenticate the presence of public figures, it also introduced a hierarchy among users—one that arguably amplified the influence of verified accounts in online debates [48].
Despite these concerns, in November 2022, following Elon Musk’s acquisition of Twitter, verification became a paid feature available through X’s Premium subscription [24].
We conduct our analysis around three key political events in Italy during 2022: the election of the President of the Republic, the political crisis, and the subsequent national elections. Our initial observation is that a strong signal of political alignment is detectable in the data: when clustering verified politicians in the retweet network by political party, the resulting partition shows high modularity. We then explore the limitations of existing community detection algorithms. When comparing their output to manually annotated data, both descriptive and inferential methods often produce noisy and imprecise community structures.
To address this, we evaluate the performance of two algorithms designed to identify discursive communities: $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, both of which begin by labeling verified users, given their central role in public debate. The $\mathrm{MonoDC}_{V}$ approach is grounded in a typical political dynamic, where less prominent figures tend to retweet content from party leaders. Conversely, the $\mathrm{BiDC}_{V}$ method—originally proposed by [18]—builds an undirected similarity network between content creators by analyzing audience overlap. In both cases, we reduce the effect of random noise—an inherent feature of OSNs [28, 29]—by comparing empirical data to a maximum entropy null model, which accounts for user activity [30].
Both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ successfully reproduce the organization of political debates into discursive communities shaped by parties or alliances. Notably, using verified users as leading content creators consistently outperforms alternative selections based solely on user activity. This suggests that the status conferred by verification plays a more decisive role in structuring the online conversation than activity alone.
In summary, our work offers two take-home messages. First, in X/Twitter datasets from before November 2022, leveraging verified users allows for a clearer and more accurate identification of discursive communities. Alternative approaches relying solely on user activity result in noisier partitions, suggesting that user status is a major factor in shaping the online debate.
Second, our findings are based on comparisons conducted before account verification became a paid feature. This raises concerns about the potential consequences of monetized verification: by enabling users to purchase visibility and prestige, the platform may open the door to malicious actors seeking to manipulate discourse and distort the public agenda. A deeper comparison between “old” –i.e. pre-November 2022– and “new” –i.e. post-November 2022– verified users will be the target of further research.
Competing interests
The authors declare that they have no competing interests.
Acknowledgements
The authors are thankful to Lorenzo Federico for interesting discussions. SG and FS were partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 “Education and Research” - Component C2 - Investment 1.1 - Next Generation EU “Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9.
Appendix A Maximum entropy null models for bipartite networks
A.1 Bipartite Configuration Model
The framework described in Section 2.1 can be adapted to bipartite networks. In particular, in the following we will use the degree sequences of nodes of both layers as constraints: in this case, the null-model takes the name of Bipartite Configuration Model (or BiCM [40]).
Let us call the two layers of a bipartite network $\top$ and $\bot$ and let $N_{\top}$ and $N_{\bot}$ be their dimensions; nodes belonging to the $\top$-⊤ -layer will be denoted by Latin indices, while the ones belonging to the $\bot$-⊥ -layer will be denoted by Greek ones. Any bipartite network is completely described by its biadjacency matrix $\mathbf{B}$, i.e. a $N_{\top}\times N_{\bot}$ matrix whose entries $b_{i\alpha}=1$ if nodes $i\in\top$ is connected to $\alpha\in\bot$ and $b_{i\alpha}=0$ otherwise. The degree of a node $i\in\top$ ($\alpha\in\bot$), i.e. the number of connections that the nodes have, is simply $k_{i}=\sum_{\alpha}b_{i\alpha}$ ($h_{\alpha}=\sum_{i}b_{i\alpha}$).
Following the procedure sketched in Section 2.1 returns that the probability of observing the generic (bipartite) graph $G_{\text{Bi}}$ is
$\begin{split}P(G_{\text{Bi}})=&\prod_{i,\alpha}\dfrac{e^{-(\theta_{i}+\eta_{\alpha})b_{i\alpha}(G_{\text{Bi}})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}\end{split}$ (3)
where $\vec{\theta}$ and $\vec{\eta}$ are the Lagrangian multipliers associated with the degree sequences on, respectively, layers $\top$ and $\bot$. Eq. (3) indicates that probabilities per graph factorize in terms of independent probabilities per link $p_{i\alpha}=\frac{e^{-(\theta_{i}+\eta_{\alpha})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}$. To get the numerical values of the Lagrangian multipliers, we can impose the likelihood maximization [38].
In the present analysis, we used the python module bicm to calculate the numerical values of the probabilities: bicm, described in Ref. [49] as a submodule of NEMtropy, is also a standalone module.
A.1.1 BiCM validation of the monopartite projection of a bipartite network
Using BiCM as a benchmark permits to validate the co-occurrences in the real network [50]. In fact, observed co-occurrences are simply the “projection” of the information contained in the bipartite network into one of the two layers, say $\top$:
$V^{ij}=\sum_{\alpha}b_{i\alpha}b_{j\alpha},\,\forall i\neq j\in\top$
(in Ref. [40] co-occurrences were also called V-motifs since, if the two layers of the bipartite networks are drawn horizontally, they sketch a “V” between the layers). Since in BiCM the probabilities per link are independent, then
$\langle V^{ij}\rangle=\sum_{\alpha}p_{i\alpha}p_{j\alpha},\,\forall i\neq j\in\top.$
In practice, the entire BiCM distribution of each $V^{ij}$ can be calculated: they are Poisson-Binomial distributions, i.e. the extension of Binomial distribution in which each event has a different probability. Calculating exactly the Poisson-Binomial distribution is computationally costly. It was shown, that when probabilities are relatively small –as in our case–, the Poisson-Binomial distribution can be finely approximated with a Poisson distribution [51]. Therefore, it is possible to efficently calculate the p-value associated with each observed co-occurrence $(V^{ij})^{*}$. Since the number of different co-occurrences is $\binom{N_{\top}}{2}$, we need multiple hypothesis testing. In the present manuscript, we implemented the False Discovery Rate (FDR,  [46]) since it controls the rate of False Positives. Again, also the validated projection used in the present manuscript was calculated using the python module bicm.
A.2 Bipartite Directed Configuration Model
Let us now consider the case in which we have a bipartite directed network and we want our model to include the information about the in- and out-degrees of nodes on both layers. It can be shown that the probability per bipartite directed graph $G_{\text{BiD}}$ factorises in terms of the BiCM probability per two undirected bipartite graphs $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, describing, respectively, all edges from layer $\top$ to layer $\bot$, and vice versa [41, 18]:
$P_{\text{BiDCM}}(G_{\text{BiD}})=P_{\text{BiCM}}(G_{\text{Bi}}^{\top\rightarrow\bot})\cdot P_{\text{BiCM}}(G_{\text{Bi}}^{\top\leftarrow\bot}).$
If $b_{i\alpha}^{\top\rightarrow\bot}$ and $b_{i\alpha}^{\top\leftarrow\bot}$ are the generic entries of the biadjacency matrices associated to, respectively, $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, then:
$\begin{split}P_{\text{BiDCM}}(G_{\text{BiD}})=&\Big{[}\prod_{i,\alpha}(q_{i\alpha}^{\top\rightarrow\bot})^{b_{i\alpha}^{\top\rightarrow\bot}}(1-q_{i\alpha}^{\top\rightarrow\bot})^{(1-b_{i\alpha}^{\top\rightarrow\bot}})\Big{]}\\ &\cdot\Big{[}\prod_{i^{\prime},\alpha^{\prime}}(q_{i^{\prime}\alpha^{\prime}}^{\top\leftarrow\bot})^{b_{i\alpha}^{\top\leftarrow\bot}}(1-q_{i\alpha}^{\top\leftarrow\bot})^{(1-b_{i\alpha}^{\top\leftarrow\bot}})\Big{]},\end{split}$
where $q_{i\alpha}^{\top\rightarrow\bot}$ and $q_{i\alpha}^{\top\leftarrow\bot}$ are the (independent) probability to observe a link from $i$ to $\alpha$ or vice versa. Following the same procedure described above for BiCM, $q_{i\alpha}$s can be expressed in terms of the Lagrangian multipliers associated with the (constrained) maximization of the entropy, i.e.
$\begin{split}q_{i\alpha}^{\top\rightarrow\bot}=&\dfrac{e^{-(\theta_{i}^{\text{out}}+\eta_{\alpha}^{\text{in}})}}{1+e^{-(\theta_{i}^{\text{out}}+\eta_{\alpha}^{\text{in}})}};\\ q_{i\alpha}^{\top\leftarrow\bot}=&\dfrac{e^{-(\theta_{i}^{\text{in}}+\eta_{\alpha}^{\text{out}})}}{1+e^{-(\theta_{i}^{\text{in}}+\eta_{\alpha}^{\text{out}})}}.\end{split}$
The numerical values of the $\theta_{i}^{\text{out}}$, $\theta_{i}^{\text{in}}$, $\eta_{\alpha}^{\text{out}}$ and $\eta_{\alpha}^{\text{in}}$ can be obtained through the maximization of the likelihood [38].
A.2.1 BiDCM validation of the monopartite projection of a bipartite network
Following the procedure described in Appendix A.1.1, we can provide a statistical validation of the monopartite multilink directed network obtained by projecting, say, on layer $\top$, a bipartite directed network. The number of edges from node $i$ to node $j$ in the projection is
$\mathcal{V}^{ij}=\sum_{\alpha}b_{i\alpha}^{\top\rightarrow\bot}b_{j\alpha}^{\top\leftarrow\bot}.$
The probability distribution of $\mathcal{V}^{ij}$, following the BiDCM, is again a Poisson-Binomial and, since also in the present case the probabilities are small, the Poisson-Binomial can be finely approximated by a Poisson distribution. Therefore we are interested in calculating $\langle\mathcal{V}^{ij}\rangle$ because it is the only parameter of the Poisson distribution. Since all $b_{i\alpha}^{\top\rightarrow\bot}$ and $b_{i\alpha}^{\top\leftarrow\bot}$ are independent, it is simply
$\langle\mathcal{V}^{ij}\rangle=\sum_{\alpha}q_{i\alpha}^{\top\rightarrow\bot}q_{j\alpha}^{\top\leftarrow\bot}.$ (4)
A.2.2 The application of BiDCM-validated projection to the bipartite network of users and posts
In the analyses above, the BiDCM-validated projection was used to validate the multilink retweet network among users. The original system is a bipartite directed network in which the two layers $\top$ and $\bot$ are, respectively, users and tweets: an arrow goes from a user to a tweet if the user is the author of the tweet and vice versa if the user retweeted the tweet. Since every post has a single author, in $G_{\text{Bi}}^{\top\rightarrow\bot}$ each node $\alpha\in\bot$ has $\kappa_{\alpha}^{\text{in}}=\sum_{i}b_{i\alpha}^{\top\rightarrow\bot}=1$. With such an identification, the probability that the user $i$ is the author of the post $\alpha$ is independent of $\alpha$ and reads:
$q_{i\alpha}^{\top\rightarrow\bot}=\dfrac{e^{-(\theta_{i}^{\text{out}}+\eta^{\text{in}})}}{1+e^{-(\theta_{i}^{\text{out}}+\eta^{\text{in}})}};$
in this case, the Bipartite Configuration Model reduces to a Bipartite Partial Configuration Model [50]. Imposing that the constraints, i.e. the number of tweets authored by each user and the fact that each post has a single author, returns
$q_{i\alpha}^{\top\rightarrow\bot}=\dfrac{\kappa_{i}^{\text{out}}}{N_{\bot}},\quad\forall\alpha\in\bot.$ (5)
Such identification is particularly effective in our case, since plugging Eq. (5) in Eq. (4), gives
$\langle\mathcal{V}^{ij}\rangle=\sum_{\alpha}q_{i\alpha}^{\top\rightarrow\bot}q_{j\alpha}^{\top\leftarrow\bot}=\dfrac{\kappa_{i}^{\text{out}}}{N_{\bot}}\sum_{\alpha}q_{j\alpha}^{\top\leftarrow\bot}=\dfrac{\kappa_{i}^{\text{out}}\kappa_{j}^{\text{in}}}{N_{\bot}},$ (6)
where we used that the degree sequence is constrained in $G_{\text{Bi}}^{\top\leftarrow\bot}$. The result in Eq. (6) is remarkable: the expected value $\langle\mathcal{V}^{ij}\rangle$ can be easily calculated without knowing explicitly the value of the Lagrangian multipliers. Therefore, p-values against BiDCM of the observed $\mathcal{V}^{ij}$ can be easily calculated, and further statistically validated using the FDR procedure described above [46].
A.2.3 Limits of applicability
Besides the simplicity of Eq. (6), there is an important issue, already mentioned for the present application in Ref. [18]. In the null model implemented above, constraints are satisfied on average. Therefore, there are configurations in the ensemble that are not physical, i.e. posts that may have either more than a single author or none. Since both non-physical configurations, i.e. more authors or none, contribute to calculating the p-values in opposite directions, their final effect is hard to predict. Nevertheless, in light of the present results and the ones of previous research, the agreement with the ground truth seems to indicate that, at least in the present application, the final effect of such an approximation is limited. Such an issue is going to be the target of further research.
Appendix B The V-measure for clustering similarity
Let us define, respectively the homogeneity and the completeness of $Q$ with respect to $P$ as
$h(P,Q)$ $=1-\frac{S(P\mid Q)}{S(P)}$
$c(P,Q)$ $=1-\frac{S(Q\mid P)}{S(Q)}.$
Note that $h=1$ if and only if the clusters obtained with $Q$ are sub-clusters of those obtained with $P$, whereas $c=1$ if and only if the clusters obtained with $Q$ are super-clusters of those obtained with $P$. It can be shown that the V-measure $\mathrm{VM}_{\beta}(P,Q)$ is the harmonic mean of $h$ and $c$, weighted by the parameter $\beta$ [31]:
$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)h(P,Q)c(P,Q)}{\beta h(P,Q)+c(P,Q)}.$
It is worth noting that the VM of two partitions does not decrease linearly with the number of elements on which the two partitions disagree: the $\mathrm{VM}_{1}$ between two partitions obtained with the same algorithm is generally not greater than 0.9; if we consider two partitions that differ by only a small percentage of the network that is randomly relabeled, their $\mathrm{VM}_{1}$ is about 0.8 with $\sim$5of the network relabeled, and decreases to about 0.6 when $\sim$10-13of the network has been relabeled (cfr. Fig. 9 in Appendix B). Based on our tests, we deduce that a reasonably good $\mathrm{VM}_{1}$ between two partitions obtained with different methods/algorithms lies in the range $[0.6-0.9]$.
Fig. 9 shows the average $\mathrm{VM}_{1}$ between some partition of the Elections dataset and a noisy version of the same partition obtained randomly relabeling part of the users. We consider two choices for the initial labeling: 5 equally sized random communities and the communities obtained running the Louvain algorithm. In both cases, the $\mathrm{VM}_{1}$ rapidly decreases to about 0.8 when $\sim$5of the network has been relabeled, and to 0.6 when $\sim$10-13of the network has been relabeled. We also show that two different runs of the Louvain algorithm generate partitions whose $\mathrm{VM}_{1}$ is about 0.9, on average.
Figure 9: Understanding the values of $\mathrm{VM}_{1}$. The plot shows $\mathrm{VM}_{1}$ between independent runs of the Louvain algorithm compared with the $\mathrm{VM}_{1}$ between some partition of the Elections dataset –either 5 equally sized random communities, or the communities obtained running the Louvain algorithm– and a noisy version of the same partition obtained randomly reallocating part of the users, as a function of the percentage of the network that is reallocated.
Appendix C Additional analyses
In the following, we report additional figures and tables that have been left out of the main sections. For the sake of readability and clarity, all discussions can be found in the respective captions.
Figure 10: The average $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms and 100 runs of the Louvain algorithm (in green, the self-similarity of Louvain partitions). Louvain was chosen as the benchmark, but all standard methods display comparable performances. Standard algorithms show great variability in the communities assigned to verified users.
Figure 11: The $\mathrm{VM}_{1}$ between partitions of the entire network, obtained with different methods, and the propagated annotations. Let us note that, when comparing the results of the propagated labels of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}$ with the annotations of verified users propagated using the same label propagation algorithm, our methods have an advantage. Nonetheless, the performances of standard methods are worse, even when limited to annotated users.
Table 3: Mean and standard deviation of the execution time, expressed in seconds, of the different steps of $\mathrm{BiDC}_{V}$ on the Elections dataset, the largest dataset considered in this paper. These timings highlight that the cost of $\mathrm{BiDC}_{V}$ is dominated by the label propagation algorithm presented in Ref. [42].
mean std
algorithm step
Label propagation on rt network 943.76 62.14
Louvain on top network 2.64 0.03
validated projection p=0.01 69.74 2.42
Figure 12: An analysis of the role of validation ($\alpha$ denotes the significance level) in determining the partitions obtained for the $\top$ layer of $\mathrm{BiDC}_{V}$. Top: the $\mathrm{VM}_{1}$ between partitions of annotated $V$ users, obtained for different validation levels, and the annotations. Bottom: the $\mathrm{VM}_{1}$ between partitions of $V$ users, obtained for different validation levels, and those obtained with $\alpha=0.01$. Each point is a different run of the method indicated by the color and marker. The results on the non-validated network are very poor, whereas the significance level has a limited impact on the inferred partitions.
Figure 13: The distribution of the H-index in the three datasets. The threshold $H=3$ used to tell apart $\top$ and $\bot$ users is justified by the empirical H-index distribution having an elbow at 3; cumulatively, only about $1-2\%$ of the users pass the threshold.
Figure 14: Cross-comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$. The $\mathrm{VM}_{1}$ between partitions obtained with different choices of $\top$ and those obtained with $\top=V$, used as a reference, for the Elections dataset. The columns refer to the user set upon which the VM is computed. For instance, “user set = influential” means that in the second column we only consider how influential users have been labeled by different methods, temporarily ignoring the rest of the network.

### v3

@@
-
+Leveraging Content Producer Networks and User Perception to Detect Online Discursive Communities
 Abstract
-In this study, we address the challenge of detecting “discursive communities” on X/Twitter by focusing on the role of verified users as the main content creators in online political debates.
-The analysis centers on three major Italian political events in 2022 – the Presidential election, a governmental crisis, and the general elections – occurring before the introduction of paid account verification.
-We propose and compare two novel methodologies, MonoDC and BiDC, which exploit, respectively, the retweet network among users and a similarity network based on shared audiences, while integrating a maximum entropy null model to filter out the inherent noise in online social networks.
-Our results demonstrate that leveraging verified users—considered as indicators of prestige and authority—leads to significantly clear community partitions that closely reflect the actual political affiliations, outperforming standard community detection algorithms applied to the entire retweet network.
-Moreover, the comparison of different methodologies and user sets suggests that the status conferred by the blue verification tick plays a dominant role in shaping online discourse, with important implications for platform governance, especially in light of the recent shift to paid verification.
-1 Introduction
-Online communities naturally emerge from users’ tendency to follow and interact with others who share similar interests, experiences, connections, or political views [1, 2, 3].
-Such communities can provide valuable support and shared knowledge [4, 5, 6]; however, they can also form “echo chambers,” where members are exposed only to a narrow range of perspectives that reinforce their existing beliefs [7, 8, 9, 10, 11, 12, 13].
-To fully understand the influence of user communities on information diffusion and opinion dynamics, it is crucial to develop effective methods for detecting these communities and capturing their key structural features.
-In the literature, several refined approaches have been proposed to enhance standard community detection methods by tailoring them to the context of online social networks.
-For example, Citraro and Rossetti [14] incorporate account attributes to strengthen the conventional definition of communities.
-Other approaches leverage the varied roles played by different accounts in online debates.
-Although Twitter/X was initially designed as a many-to-many platform, numerous studies have observed that it actually functions as a few-to-many network, where a small number of influential content creators generate the majority of posts [15, 16, 17].
-In this context, it becomes natural to focus on the leading content creators for detecting communities on Twitter/X.
-Identifying which accounts should be considered as leading content creators is, therefore, essential.
-In [4], influential users—defined as those with more followers than followees and with more mentions received than given—are identified as the primary figures in the debate.
-Similar approaches, based on user activity, have also been explored [15, 16, 17].
-Other studies place more emphasis on account prestige than on activity, considering verified users as the leading content creators [18, 19, 20].
-Verified users on X/Twitter have an interesting history: they were first introduced in June 2009 to curb the impersonation of notable persons or organizations [21, 22].
-In this early stage, the platform assumed responsibility for identifying key public figures and verifying their identities with a blue checkmark, while standard users were not part of the verification process [23].
-Efforts to broaden the pool of verified accounts through user applications were unsuccessful [21].
-Eventually, in November 2022, verification became a feature of X’s Premium accounts, available upon payment [24]—a change that has sparked concerns regarding the influence of verified users on online debates [25, 26].
-In the present study, we investigate the method proposed in [18] using three distinct datasets collected during 2022, in a period before blue ticks could be purchased via X’s Premium subscription.
-Our framework comprises: (i) identifying verified users as the leading content creators; (ii) constructing an informative network of verified users by filtering out random noise through a comparison between empirical data and an appropriate null model; (iii) detecting communities within this sub-network; and (iv) propagating the labels to all other users.
-Following the terminology introduced by [27], the resulting partition is referred to as “discursive” communities.
-A key component of our approach, in addition to leveraging verified users, is the elimination of random noise from the system by comparing empirical data against a maximum entropy statistical benchmark.
-As numerous studies have shown, removing random noise is crucial in online social networks to extract the relevant signal regarding users’ preferences and behaviors [18, 28, 29].
-In this regard, maximum entropy null models provide an ideal statistical benchmark, being maximally random apart from the constraints imposed [30].
-Figure 1: MonoDC’s pipeline.
-Starting from the data from Twitter/X, we build a bipartite directed network between accounts and posts; a directed arrow from a user to a post represents authorship, while in the opposite direction represents a retweet.
+Online discussions are often characterized by strong behavioral asymmetries: a relatively small fraction of users actively produces content, while the majority primarily consumes and redistributes it.
+Here we propose a community-detection framework for online social networks that exploits this asymmetry by first identifying and clustering a set of leading users, and then extending the resulting labels to the broader user base.
+We introduce two complementary strategies to cluster leaders, one based on their mutual interactions and the other on audience overlap, both relying on entropy-based filtering to separate signal from noise.
+We evaluate the framework on three major Italian political debates on Twitter/X, using public figures–identified through the pre-2022 verification system–as leaders, and known affiliations of political actors as ground truth labels.
+Compared with standard baselines, the proposed approach yields more coherent and interpretable communities aligned with political structures, with the two variants respectively recovering parties and coalitions.
+Activity-based criteria for selecting leaders produce qualitatively similar but consistently weaker results, particularly at the coalition level.
+Overall, our findings show that creating statistically validated networks of publicly recognized figures, whose off-platform roles constrain and stabilize their online behavior, provide a strong basis to identify discursive communities on social media.
+Although developed for Twitter/X, the approach is conceptually general, as it leverages structural asymmetries common to many online platforms.
+Introduction
+Online social platforms have become crucial arenas for public discourse, with communities forming organically around shared narratives, values, and ideological affinities [Conover2011a, Conover2011b, Conover2012].
+While these discursive communities can offer solidarity and foster public engagement [GonzalezBailon2013, Andalibi2018, Morini2025], they also risk becoming echo chambers—insulated environments that reinforce existing beliefs and suppress dissenting perspectives [jamieson08echo, Garrett2009, delvicario2016, Zollo2017Debunking, Villa2021, Morini2021, Pratelli2024].
+Understanding the structural dynamics that govern these communities is therefore essential to evaluating how information circulates and how opinions crystallize in digital spaces.
+Traditional approaches to community detection in social networks have evolved significantly, incorporating network topology, account metadata, and behavioral signals.
+For instance, some methods enhance classical algorithms by integrating account characteristics [citraro2019eva], while others acknowledge that platforms like Twitter/X function less as egalitarian many-to-many networks, and more as hierarchies dominated by a small number of highly visible content creators [Watts2007, Wu2011, Hilbert2017].
+Within this structure, influence is disproportionately concentrated, making the identification of these “elite” accounts central to the analysis of online discourse [Becatti2019d, Caldarelli2020b, Radicioni2021a].
+Building on this body of work, the first contribution of this article is the definition of a general framework for detecting online discursive communities, formalizing the idea that online debates are driven by a relatively small set of users who actively produce content.
+Because content production entails greater exposure and commitment—especially in political discussions—these users tend to adopt more consistent and publicly interpretable positions.
+Once content producers are identified, the affiliation of less active users can be inferred from their interaction patterns, even when individual behavior is sparse.
+The proposed framework operationalizes this rationale in a relatively simple way.
+We partition the user set $U$ into two disjoint subsets, $U=\top\cup\bot$, cluster the content creators (the set $\top$) by applying a standard community-detection algorithm to an appropriately defined network $G=(\top,E)$, and then extend the resulting labels to users in $\bot$ based on their interactions with $\top$.
+A key modeling choice, however, concerns the construction of the edge set $E$, which defines how relationships among content creators are represented.
+We consider two alternatives: one based on statistically significant interactions directly observed among users in $\top$, and one based on a validated projection of the bipartite network linking content creators in $\top$ to their audiences in $\bot$.
+These two variants define Monopartite-network-induced Discursive Communities ($\mathrm{MonoDC}_{\top}$) and Bipartite-network-induced Discursive Communities ($\mathrm{BiDC}_{\top}$), respectively.
+Both methods rely on maximum-entropy null models [Cimini2019] to statistically validate inferred connections and to filter out noise–a well-known challenge in the analysis of online social networks [Becatti2019d, Declerck2022a].
+Statistical filtering plays a limited role in $\mathrm{MonoDC}_{\top}$, but is essential in $\mathrm{BiDC}_{\top}$ to remove spurious correlations induced by heterogeneous audience activity.
+Beyond the construction of $E$, the framework admits different criteria to identify the set $\top$ of content creators.
+From this perspective, the second contribution of this article is less methodological and more substantive: we show that, although social-media relevance can be acquired in an anonymous and platform-driven manner, ideological communities in political debates remain predominantly anchored to publicly recognizable political figures.
+Accordingly, we focus on Twitter’s verified accounts ($\top=V$), which, at the time covered by our datasets, identified public figures whose activity on the platform was closely tied to their off-platform roles.
+All datasets analyzed precede the introduction of paid verification in November 2022, and therefore correspond to this original, institutionally grounded meaning of verification [Haman2023].
+We compare this choice with two activity-based criteria derived from the data.
+We consider influential users ($\top=I$) as defined by [GonzalezBailon2013], based on follower structure and mention activity, and high-$h$-index users ($\top=H$), obtained by adapting the Hirsch index [Hirsch2005] to retweet dynamics.
+These data-driven selections serve as baselines to evaluate the added value of anchoring the analysis on publicly recognized figures.
+We analyze three datasets covering major political events in Italy in 2022, where discursive communities are expected to mirror political alignments.
+When evaluated against the known affiliations of clearly identifiable political actors, $\mathrm{MonoDC}_{V}$ captures party-level structures, whereas $\mathrm{BiDC}_{V}$ is more effective at identifying coalitional configurations.
+Standard community-detection algorithms applied to the complete retweet network yield noisier and less interpretable partitions, while activity-based criteria perform better than these baselines but remain less reliable than anchoring the analysis on verified accounts, especially at the coalition level.
+We argue that this is not an effect of verification, but, again, a consequence of the fact that, historically, only publicly recognizable figures were “verifiable”, a requisite that in turn constrained and stabilized their online behavior.
+In this sense, our results show that, even in environments where visibility can be gained in largely anonymous and platform-driven ways, ideological communities in political debates remain primarily structured around actors endowed with public recognizability beyond the platform.
+By foregrounding this mechanism, we propose and analyze a framework that, while developed for Twitter/X, is conceptually general, as it leverages asymmetries between content producers and audiences that characterize many contemporary online platforms.
+Finally, we show that a carefully selected subset of users and interactions captures a large fraction of the information needed to reconstruct political communities, offering practical insights for designing efficient data-collection strategies under increasingly restrictive platform access conditions.
+Related work
+A substantial body of research has investigated the structure of the online discourse on Twitter and other social media platforms by analyzing retweet networks, information cascades, and statistically validated interaction patterns (e.g., [garimella2018quantifying, Becatti2019d, cinelli2021echo, lorenzspreen2023digitalmedia]).
+These studies have consistently shown, in particular, that political debates on social media exhibit strong polarization and modular structure, often aligned with ideological or partisan divides.
+A central methodological challenge in this literature concerns the high level of noise and heterogeneity in user activity.
+To address this issue, several works have adopted statistical filtering techniques based on null models to extract significant interaction patterns from bipartite or projected networks.
+In particular, maximum-entropy null models have been widely used to preserve local constraints—such as node degrees—while filtering out correlations that can be explained by random activity alone [Cimini2019, Saracco2017].
+These approaches have been shown to uncover meaningful mesoscale structures in retweet, mention, and user-content networks, enabling the identification of latent ideological alignment from interaction data [Becatti2019d, Caldarelli2020b, Radicioni2021a, Mattei2022, Pratelli2024].
+Within this line of work, the Italian political context has received special attention.
+Becatti et al.
+[Becatti2019d] have shown that entropy-based filtering improves the interpretability and stability of political communities compared to unfiltered network representations (the method proposed therein is called $\mathrm{BiDC}_{V}$ in the present manuscript).
+Related studies have further demonstrated that statistically validated user–hashtag bipartite projections can reveal well-defined discursive communities during national electoral campaigns, even in the absence of explicit ideological labels [Radicioni2021a].
+A parallel strand of research has focused on the role of influential or central users in shaping the online discourse.
+Early work has defined influence primarily in terms of activity-based or structural measures, such as the number of followers, retweets, mentions, or centrality in interaction networks [Watts2007, Hilbert2017, Wu2011].
+Other studies have explored the diffusion power of opinion leaders and elites in online debates, often highlighting the disproportionate impact of a small fraction of highly visible users [bakshy2011everyone].
+Despite these advances, two important gaps persist in the literature.
+First, most community-detection approaches implicitly treat users symmetrically, without explicitly accounting for the strong asymmetry between content producers and content consumers that characterizes participation on social media–in general, and especially for political debates.
+Second, benchmarks for evaluating communities are often based on internal network metrics, such as modularity or stability, rather than on externally validated ground truth labels.
+As a result, it remains unclear (i) whether institutional signals of authority–such as pre-2022 verified status on Twitter–provide particularly informative anchors for community detection, and (ii) how alternative criteria for identifying leading users affect the resulting mappings when embedded into statistically validated pipelines.
+The present work addresses these gaps by proposing a leader-based framework that explicitly separates content creators from their audiences and systematically compares different strategies for selecting politically relevant producers of content.
+Methods
+We investigate how online discursive communities form around publicly relevant users on Twitter/X by analyzing large-scale interaction data.
+While the framework we propose is general, we focus empirically on three major political events in Italy during 2022, a domain in which online debates tend to be especially structured and externally grounded.
+Our approach builds on the observation that social-media debates are typically characterized by strong asymmetries in user behavior: a relatively small subset of users actively produces salient content, whereas the majority primarily consumes and redistributes it.
+This asymmetry is particularly pronounced in political discussions, where public figures and institutional actors play a central role.
+To exploit this structure, we adopt a leader-based framework in which community detection is first performed on a selected set of content creators and subsequently extended to the broader user base.
+Detection of political discursive communities
+To identify meaningful structures in online political debates, it is necessary to distinguish statistically significant interaction patterns from correlations induced by heterogeneous user activity and content popularity.
+To this end, our approach relies on maximum-entropy null models as a principled statistical baseline for validating inferred network connections before applying community-detection algorithms.
+Within this framework, an ensemble of random networks is defined by maximizing the Shannon entropy subject to a set of local constraints that reproduce selected properties of the observed data, such as degree sequences or activity levels [Jaynes1957, park2004statistical, Cimini2019].
+Empirical links are evaluated against their expected occurrence under the null model, and only statistically significant deviations are retained.
+Model parameters are estimated by maximizing the likelihood of the observed network [Garlaschelli2008], and statistical significance is assessed using false discovery rate correction [Benjamini1995].
+This maximum-entropy approach is grounded in information theory and statistical mechanics and has proven effective across a wide range of networked systems [Becatti2019d, Caldarelli2020b, Caldarelli2021, Radicioni2021a, Radicioni2021b, Mattei2021, Guarino2021, Mattei2022, Bruno2022, Declerck2022a, Declerck2022b, Pratelli2024].
+In the following, we introduce two complementary methods that instantiate this framework in different ways, depending on how interactions among selected content creators are inferred.
+Further methodological details are reported in Section A of the Appendix.
+In the analyses presented here, we focus on verified accounts as a concrete instantiation of content creators, as–prior to the introduction of paid verification–they largely corresponded to publicly recognized figures.
+However, the methods described below are general and can be applied to any set $\top$ of users identified as debate leaders.
+Construction of the $\mathrm{MonoDC}_{V}$ network
+The first method, $\mathrm{MonoDC}_{V}$ (Monopartite-network-induced Discursive Communities; Fig.
+1), focuses on retweet interactions among verified users.
+We start by constructing a bipartite directed network between users and tweets, where directed edges represent authorship and retweet actions.
+At this stage, we consider all verified users together with all users who retweeted at least one verified account in the dataset under study.
+This structure is then projected onto a weighted directed user-user network, in which a link from user $i$ to user $j$ encodes the number of times $i$ retweeted content authored by $j$.
+To separate meaningful interaction patterns from noise–an intrinsic feature of online social activity–we apply a maximum-entropy null model tailored to bipartite directed networks.
+The model uses the observed posting and retweeting activity of users as constraints while randomizing all other aspects, providing a principled statistical baseline.
+For each pair of users $i,j$, we compute the probability of the observed retweeting weight under the null model, and retain the edge only if its p-value is statistically significant (more details are present in Section A of the Appendix).
+Community detection is then performed on the subnetwork of the validated network induced by verified users.
+The presented results have been obtained with a modularity-based algorithm, but we verified that alternative algorithms yield essentially the same results.
+Finally, the detected community labels are propagated to non-verified users using a standard label-propagation scheme [Raghavan2007a].
+Figure 1: $\mathrm{MonoDC}$’s pipeline.
+Starting from the data from Twitter/X, we build a bipartite directed network between accounts and posts; a directed arrow from a user (a circle) to a post (a square) represents authorship, while in the opposite direction represents a retweet.
 Then, the content of the bipartite network is projected into a monopartite directed network of users.
 Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite directed networks.
-Then, the sub-network relative to content creators is analysed, and communities are detected using standard algorithms.
-Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
-Compared to previous work, we distinguish between two main versions of the algorithm that differ in how step (ii) is implemented.
-In $\mathrm{MonoDC}$ (Monopartite-network-induced Discursive Communities), we consider the direct interactions —i.e., retweets— among content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig.
-1.
-In $\mathrm{BiDC}$ (Bipartite-network-induced Discursive Communities), we instead construct a similarity network based on the shared audience between pairs of leading content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig.
-2.
-Additionally, we explore various classes of leading content creators beyond verified users.
-Our proposed methods were tested on three large datasets linked to the primary political events in Italy during 2022, before user verification became a feature of Premium accounts.
-To validate our results, we manually annotated approximately 300 verified accounts to determine their political leanings and compared different clustering algorithms to assess their ability to detect this signal.
-The key finding is that both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$, when using verified users as the main content creators, significantly outperform standard algorithms.
-They produce clearer partitions that consist of a few meaningful communities, which accurately reflect the political landscape during the studied period.
-Specifically, $\mathrm{MonoDC}$ appears to capture political parties more effectively—likely because politicians predominantly retweet their party leaders—while $\mathrm{BiDC}$ better identifies political alliances due to the shared construction of narratives.
-Importantly, our results deteriorate markedly when the set of main content creators is altered.
-Figure 2: BiDC’s pipeline.
-Starting from the data from Twitter/X, we build a bipartite undirected network between the accounts of the leading content creators and the one of standard users; a link is present if the standard users retweeted the given content creator at least once.
-Then, the information of the bipartite network is projected into a monopartite network of leading content creators.
+Then, the sub-network relative to the selected content creators (magenta circles) is analysed, and communities are detected using standard algorithms.
+Finally, the so-obtained labels are propagated to standard users (blue circles) through a label propagation algorithm.
+The icon representing Twitter/X data is obtained by superimposing the icons “Data Base #7994662” by Farrih Icon and “Hashtag #4315467” by Alex Burte, both from Noun Projects.
+Construction of the BiDCV network
+The second method, $\mathrm{BiDC}_{V}$ (Bipartite-network-induced Discursive Communities; Fig.
+2[Becatti2019d]), infers relationships among verified users indirectly, based on the similarity of their audiences.
+Here, we construct a bipartite undirected network connecting verified users to all users who retweeted them at least once during the event.
+This bipartite structure preserves the heterogeneity between content creators and their audiences.
+To extract meaningful similarities between verified users, we apply a maximum-entropy null model for bipartite undirected networks, which accounts for differences in audience sizes and engagement patterns.
+Statistically significant similarities are retained, yielding a validated projection onto a network of verified users.
+Community detection is performed on this projected network, and the resulting labels are propagated to non-verified users according to the composition of their retweet activity.
+As in $\mathrm{MonoDC}_{V}$, the propagation step allows us to assign community memberships to the entire user set starting from a limited number of leader nodes.
+Figure 2: $\mathrm{BiDC}$’s pipeline.
+Starting from the data from Twitter/X, we build a bipartite undirected network between the accounts of the content creators and the one of standard users; a link is present if the standard users retweeted the given content creator at least once.
+Then, the information of the bipartite network is projected into a monopartite network of content creators.
 Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite undirected networks.
 Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
-In summary, our study demonstrates that verified accounts play a central role in shaping discursive communities on X/Twitter.
-This evidence carries important implications, especially considering that our datasets are limited to the period before the implementation of X’s Premium accounts, when blue checkmarks were granted based on the platform’s autonomous verification of key accounts.
-While we cannot conclusively determine whether the effectiveness of verified users as leading content creators stems solely from the prestige of the verification or from other factors, our findings indicate that definitions based exclusively on user activity yield inferior performance.
-If the prominence of verified users in steering the online discourse is indeed due to the prestige associated with blue ticks, then introducing verification as a paid feature could pose serious challenges to the platform’s integrity, potentially enabling malevolent actors to purchase authority to distort the political agenda.
-The paper is organised as follows.
-In Section 2, we introduce the methodology adopted in this study.
-Specifically, we present maximum entropy null models and their application to the detection of discursive communities (Subsections 2.1 and 2.2).
-We also define the V-measure [31], the metric used to compare different partitions (Subsection 2.3).
-Section 3 presents our results.
-We begin by illustrating the limitations of standard methods in reproducing non-trivial benchmarks observed in real-world data (Subsection 3.3).
-We then show how $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ address these issues when verified users are identified as the primary content creators (Subsection 3.4).
-Additionally, we show that the performance of both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ deteriorates when different groups of leading content creators are considered (Subsection 3.5).
-Finally, in Section 4, we summarise our findings and discuss their potential implications.
-2 Methods
-2.1 Maximum-entropy null models and discursive communities
-As underlined by many studies across a multitude of applications  [18, 19, 20, 27, 32, 33, 34, 35, 36, 28, 29, 13], using a proper statistical validation is critical for the analysis of online social networks (OSNs) that are particularly prone to noise.
-Entropy-based null models [30] represent a tailored and, at the same time, unbiased tool for detecting the statistically significant signal in complex networks.
-The method essentially consists in: i) defining an ensemble of graphs, i.e.
-the set of all possible graphs having the same number of nodes as the observed network; ii) performing a constrained maximization of the Shannon entropy of the ensemble, fixing some constitutive properties of the original system [37]; iii) maximizing the (log-)likelihood of the observed network [38], to get the numerical values of the parameters of an effective fitness model.Remarkably, this framework mimics Jaynes’ derivation of statistical ensembles in Physics starting from Information Theory [39].
-Let us start from a real network $G^{*}$ (in the following, all quantities related to the observed network will be denoted by an asterisk *).
+The procedure was first proposed in [Becatti2019d].
+The icon representing Twitter/X data is obtained by superimposing the icons “Data Base #7994662” by Farrih Icon and “Hashtag #4315467” by Alex Burte, both from Noun Projects.
+Evaluation
+We evaluate the quality of the detected communities using the V-measure ($\mathrm{VM}_{\beta}$[rosenberg2007v], an information-theoretic metric that quantifies the agreement between two partitions based on their mutual information.
+The V-measure is well suited to our setting, as it remains interpretable across different levels of partition granularity and explicitly balances homogeneity–how well each cluster contains only members of a single class–and completeness–whether all members of a given class are assigned to the same cluster–through the parameter $\beta$.
+For $\beta=1$, $\mathrm{VM}_{1}$ reduces to the Variation of Information [Meila2007], normalized by the total entropy of the partitions.
+By varying $\beta$ over $0,+\infty)$, we account for both finer-grained and more aggregated political structures, such as political movements/trends within parties or political coalitions.
+More details can be found in Section B of the Appendix.
+As a baseline, we also apply standard community-detection algorithms directly to the complete retweet network.
+This network is defined as a directed, weighted graph including all active users, where edges represent the number of retweets between pairs of users.
+Unlike $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, this representation is not statistically filtered and serves to isolate the contribution of entropy-based validation and leader selection.
+Although MonoDCV and BiDCV ultimately rely on Louvain for final partitioning, the comparison with Louvain applied to the complete retweet network serves a different purpose.
+In our frameworks, Louvain operates after statistically validated filtering, and with an informative initialization provided by verified seeds.
+Running Louvain directly on the raw network therefore represents a baseline that isolates the contribution of our preprocessing and seed selection.
+This allows us to quantify the extent to which discursive communities emerge from the raw interaction structure alone versus from statistically validated and leader-anchored representations.
+Data
+We analyze three datasets covering the Italian presidential election, a government crisis, and the 2022 general elections.
+Each dataset consists of millions of tweets and hundreds of thousands of users, collected via Twitter’s Academic Application Programming Interface (API) using event-specific keywords; more details about the queries used for the data collection can be found in Section C of the Appendix..
+Table 4 summarizes the main characteristics of the datasets.
+Among the verified accounts appearing in each dataset, we manually annotated those corresponding to politically active public figures.
+Annotations were performed at both the party and coalition level and provide an external ground truth against which we assess the performance of our methods.
+Manual annotation of verified accounts
+For each event, we manually annotated the political affiliation of verified users to construct the ground truth used in our evaluation.
+Only verified accounts of political parties and politicians were annotated; accounts belonging to journalists, media outlets, institutions, or entertainers were excluded.
+Each annotated user was assigned both a party label and a coalition label.
+The party label was determined based on publicly available information from the user’s Twitter profile and Wikipedia page, or from institutional sources such as party websites, parliamentary records and candidate lists.
+The coalition label was assigned according to the official coalitions recognized during the 2022 Italian electoral cycle.
+Annotations were performed by one of the authors, while the remaining authors conducted random spot checks to assess consistency and correctness.
+Verified accounts with no clear political alignment were included in the construction of both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ networks, but were excluded from the ground truth and used only as unlabelled nodes in the evaluation.
+Table 1: A summary of the datasets considered in the present paper
+[…]
+President1 119.018 711.497 292 1148 5804 1296
+Crisis2 113.876 1.414.181 229 1004 7038 2554
+Elections3 222.356 2.784.951 239 1555 12184 4382
+  • 1
+    Debate around the election of the President of the Republic. The political debate ended with the re-election of Sergio Mattarella on January 29, 2022.
+  • 2
+    Debate around the political crisis that led to the fall of the Draghi government on July 21, 2022.
+  • 3
+    Debate around the Italian general elections held on September 25, 2022, that led to Giorgia Meloni becoming the new Prime Minister.
+Code and data availability
+Preprocessed data and code used for the analysis are available on GitLab at https://gitlab.com/s.guarino/dico_analysis.
+Results
+Verified accounts provide a strong political signal under the pre-2022 verification regime
+To assess the effectiveness of our framework, we focused on three political events that sparked intense debate on Italian Twitter in 2022: the re-election of the President of the Republic, the resignation of Prime Minister Draghi, and the general elections in September.
+Crucially, these datasets were compiled before the platform introduced paid verification, ensuring that verified accounts retained their original role as signals of institutional recognition rather than self-declared status.
+We manually annotated a subset of verified accounts (between 229 and 292 per event) based on their publicly identifiable political affiliations.
+When restricting the retweet networks to, first, the Greatest Connected Component (GCC), and then to annotated users and evaluating the partition induced by their known party and coalition labels, we observed highly modular structures.
+Modularity values exceeded 0.6 in all cases (cfr.
+Table 2), indicating that retweet interactions among verified users are strongly structured along political lines.
+These results show that, even without any statistical validation, retweet interactions among the annotated verified users already separate clearly along party and coalition lines, indicating that well-defined political communities exist within this subset.
+Table 2: Retweet networks of annotated users and modularity (Q) of the partition induced by the manual annotation at the party and coalition level.
+Dataset Nodes in GCC Links $Q(\text{parties})$ $Q(\text{coalitions})$
+President 169 219 0.763 0.618
+Crisis 145 246 0.721 0.641
+Elections 154 355 0.777 0.649
+To extrapolate this benchmark to the broader network, we propagated community labels from verified users to the rest of the user population using a label propagation algorithm [Raghavan2007a].
+While these propagated labels are inherently noisier than manual annotations, they provide a principled proxy for large-scale political alignment and enable a systematic comparison of community-detection strategies over the full dataset.
+Standard community detection on the full retweet network shows limited political alignment
+We applied four widely used community detection algorithms—Louvain, label propagation, Infomap, and stochastic block model inference—to the complete retweet networks, defined as the directed weighted graph including all users active during the event and all observed retweet interactions.
+While all methods identified non-trivial community structure, their agreement with the manual annotations was limited.
+V-measure scores for party affiliation ranged between 0.4 and 0.6 (cfr.
+Fig.
+3), with substantial variability across algorithm runs.
+Comparisons between algorithms further revealed significant inconsistency, with pairwise V-measures for the same dataset lying in the $[0.3,0.7]$ range (cfr.
+Fig.
+12 in the Appendix).
+Qualitative inspection confirmed that standard methods often conflated ideologically distinct actors or fragmented coherent political blocs.
+These results do not indicate a general failure of standard algorithms, but rather highlight the difficulty of recovering politically meaningful communities when they are applied directly to noisy retweet networks without distinguishing between heterogeneous user roles or filtering statistically expected interactions.
+Figure 3: Limits of standard methods.
+The $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms, restricted to the annotated users, and the parties’ annotations.
+Each violin represents the distribution over 100 independent runs.
+Standard methods show a limited accuracy in recognizing the affiliation of renowned politicians.
+Leader-anchored and statistically validated networks improve community detection
+Applying our two proposed methods, $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, resulted in a substantial improvement in community-detection performance.
+When verified users were used as anchors within statistically validated networks, both methods produced partitions closely aligned with the annotated political affiliations.
+$\mathrm{MonoDC}_{V}$ achieved particularly high accuracy at the party level, reflecting the focused retweet patterns through which politicians tend to amplify content from their own party leadership.
+By contrast, $\mathrm{BiDC}_{V}$ more effectively captured coalition-level structures, as it relies on shared audience behaviour rather than direct interactions among leaders.
+Quantitatively, V-measure scores for both methods systematically exceeded those of standard algorithms across all datasets.
+$\mathrm{MonoDC}_{V}$ reached values above 0.75 for party-level annotations, while $\mathrm{BiDC}_{V}$ achieved comparably strong results at the coalition level (cfr.
+Fig.
+4).
+These findings remained robust across different values of the resolution parameter $\beta$ (cfr.
+Fig.
+5) and when considering propagated labels on the full network (cfr.
+Fig.
+13 in the Appendix).
+As a baseline, we apply the Louvain algorithm to the complete retweet network.
+This representation includes neither statistical filtering nor a distinction between verified and non-verified users, and thus reflects the raw interaction structure dominated by highly heterogeneous activity patterns.
+Although $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ also rely on Louvain for the final partitioning step, the comparison serves a different purpose.
+In our framework, Louvain is applied after maximum-entropy filtering and on networks anchored to a coherent set of political leaders.
+Running Louvain directly on the raw retweet network therefore isolates the contribution of preprocessing and leader selection, rather than acting as a competing methodological alternative.
+Notably, the Louvain algorithm performs no worse than the others in recovering the annotations (cfr.
+Fig.
+3).
+Overall, these results indicate that verified accounts function as effective anchors for political discursive communities within a statistically validated framework.
+The improved performance does not stem from verified status alone, but from its combination with maximum-entropy filtering and with the institutional coherence of the verified-user set under the pre-2022 verification regime.
+Figure 4: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods.
+The $\mathrm{VM}_{1}$ between partitions of the annotated users, obtained with different methods, and the annotations at party (first row) and coalition (second row) level.
+Each violin represents the distribution over 100 independent runs, and we use the Louvain algorithm as a benchmark because it performs no worse than other algorithms (cfr.
+Fig.
+3).
+$\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions.
+Figure 5: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods at different resolution levels.
+The $\mathrm{VM}_{\beta}$ between partitions of the annotated users, obtained with different methods, and the annotations, for different values of $\beta$ and both at the parties (top) and coalitions (bottom) level.
+Each point is the average over 100 runs of the method indicated by the colour and marker.
+Analogous results are obtained when considering the propagated annotations.
+When compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e.
+at essentially all scales at which the political spectrum can be observed.
+In this sense, standard algorithms, when executed on the entire retweet network, do not return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect.
+MonoDC captures party structure, while BiDC reveals coalitional organization
+The complementary strengths of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ arise from the nature of their respective network constructions.
+$\mathrm{MonoDC}_{V}$ relies on direct retweet interactions among verified users, which tend to encode endorsement and intra-party signalling.
+As a result, its statistically validated networks retain a large fraction of edges (94–97%), indicating dense and coherent communication among political actors.
+$\mathrm{BiDC}_{V}$, by contrast, infers similarity among leaders through audience overlap.
+After statistical validation, only 7–12% of edges are retained, reflecting the fact that only a small subset of shared audiences carries meaningful political alignment.
+This filtering allows $\mathrm{BiDC}_{V}$ to surface stable coalition-level structures, particularly in phases where electoral alliances are clearly defined.
+Consistently, the $\mathrm{VM}_{\beta}$ between the partitions produced by the two methods shows that $\mathrm{MonoDC}_{V}$ communities tend to be subclusters of those identified by $\mathrm{BiDC}_{V}$, especially for the Crisis and Elections datasets (cfr.
+Fig.
+6).
+Figure 6: Different resolution of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$.
+The $\mathrm{VM}_{\beta}$, for variable $\beta$, between partitions of all users obtained running $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$.
+The clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
+A qualitative comparison further illustrates these differences (cfr.
+Fig.
+7).
+$\mathrm{MonoDC}_{V}$ yields fine-grained partitions aligned with individual parties, whereas $\mathrm{BiDC}_{V}$ aggregates these into broader coalitional structures.
+In contrast, Louvain applied to the full retweet network produces a noisier mapping that conflates party- and coalition-level affiliations.
+We attribute this fragmentation not to political complexity, but to the structural noise of the raw retweet network, which is substantially reduced by maximum-entropy filtering.
+(a) Louvain on the retweet network.
+(b) $\mathrm{MonoDC}_{V}$.
+(c) $\mathrm{BiDC}_{V}$.
+Figure 7: Qualitative comparison of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with Louvain on the Elections dataset.
+For each of the considered methods, the Sankey diagrams show the composition of the extracted communities in terms of political parties.
+$\mathrm{MonoDC}_{V}$ very accurately detects parties (Azione and Italia Viva ran together as a single list at the elections), plus some political movements/trends within parties; $\mathrm{BiDC}_{V}$ almost perfectly detects the official political coalitions of those elections; Louvain, executed on the entire retweet network, provides a more noisy classification that is a mix of the two levels, with Forza Italia separated from its coalitions and many politicians contributing to communities led by some of their rivals.
+Activity-based definitions of leaders yield weaker political partitions
+[…]
+In all cases, performance deteriorated relative to anchoring communities on verified users (cfr.
+Fig.
+8).
+While activity-based criteria identify highly active accounts, their discursive behavior is more heterogeneous and less predictive of political alignment.
+In particular, H-index-based selections include journalists and commentators whose audiences span ideological boundaries, blurring community structure.
+These results highlight that the effectiveness of our framework depends not only on selecting active users, but on identifying content creators whose publicly recognized roles constrain and stabilize their online behavior.
+Further comparison can be found in Section D of the Appendix.
+(a) $\mathrm{MonoDC}_{\top}$, for different choices of $\top$, compared with annotations at the parties level.
+(b) $\mathrm{BiDC}_{\top}$, for different choices of $\top$, compared with annotations at the coalitions level.
+Figure 8: Comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$.
+The $\mathrm{VM}_{1}$ between partitions obtained with different choices of $\top$ and the annotations, for the Elections dataset.
+For each method, we used the level of annotation that the method more accurately reproduces, that is, parties for $\mathrm{MonoDC}_{\top}$ (a) and coalitions for $\mathrm{BiDC}_{\top}$ (b).
+The columns refer to the user set upon which the VM is computed.
+For instance, “user set = influential” means that in the second column we only consider how influential users have been labeled by different methods, temporarily ignoring the rest of the network.
+Discussion
+Digital platforms play an increasingly influential role in shaping the structure and visibility of public discourse.
+One crucial contribution of this work is the characterization of a general methodology for extracting meaningful discursive communities from noisy social data.
+We introduced a leader-based framework that explicitly accounts for the asymmetric structure of online participation.
+By separating content producers from their audiences, our approach allows communities to emerge more clearly than methods that treat all users symmetrically.
+We showed that identifying and positioning content producers is a necessary first step, after which the broader user base can be reliably assigned.
+[…]
+In addition, through a systematic comparison of three alternative criteria for selecting leading users, our findings demonstrate that institutional status, rather than raw activity, provides the strongest signal for organizing online debates.
+We found, in fact, that pre-2022 verified accounts represent the most informative anchors for community detection.
+We interpret this result as a consequence of their public recognizability and the behavioral constraints imposed by their off-platform roles, rather than as an intrinsic property of verification itself.
+While the institutional meaning of verification has changed following the introduction of paid subscriptions, the methodological insights of this study remain valid.
+In particular, our results suggest that carefully curated sets of publicly relevant actors–identified either through institutional roles or external information–can serve as effective seeds for analyzing online discourse, at least in the well-studied and highly structured context of political debates.
+This insight opens the door to more efficient data-collection strategies, in which the ego networks of a limited number of high-relevance users are sufficient to recover much of the political structure of online debates, even under restrictive platform access conditions.
+Indeed, similar strategies have been implemented, for instance in [Brugnoli2024].
+More broadly, our study reinforces the idea that the organization of digital public debate potentially depends not only on individual user behavior, but also on how platforms encode and signal authority.
+While the methods proposed here are technical, they contribute to a growing body of work showing that discursive structure, legitimacy, and visibility are deeply intertwined in online public spheres.
+As any other product of research, our analyses have some limitations.
+First, our empirical results pertain exclusively to the pre-2022 verification regime, under which verified accounts corresponded to publicly recognized figures and formed coherent, politically aligned communities.
+In this historical setting, verification acted as a reliable proxy for public relevance and institutional embeddedness, providing a stable anchor for discursive structures.
+The transition to paid verification introduces a qualitatively different regime, whose effects fall outside the scope of the present analysis.
+Concerns about potential distortions of online discourse–such as increased visibility for actors lacking public accountability or the erosion of previously coherent community structures–should therefore be interpreted as speculative and grounded in plausible mechanisms rather than in direct empirical evidence.
+Whether political discourse under the new regime will remain similarly structured, fragment into less interpretable patterns, or reorganize around alternative signals of authority remains an open empirical question.
+Future research should explicitly address this gap by analyzing data collected after the introduction of paid verification, enabling direct comparisons between the two regimes.
+Such studies could assess whether the structural role once played by verified users has weakened, persisted, or been replaced by other indicators of public relevance.
+Second, our empirical analysis focuses on political debates within a single national context, Italy, and on a single social media platform, Twitter/X.
+This choice reflects both data availability and methodological considerations: analyzing a political system that is well known to the authors allows for more reliable manual annotations and a more accurate interpretation of political actors, coalitions, and discursive dynamics.
+Importantly, this focus does not represent an intrinsic limitation of the proposed framework, which is grounded in structural features of online discourse–such as strong participation asymmetries and the presence of discussion leaders–that have been widely documented across platforms, languages, and cultural settings [Watts2007, Hilbert2017, Bracciale2018].
+Consistently with this perspective, previous studies have already shown that $\mathrm{BiDC}$ performs well in different countries [Mattei2022, Declerck2022a, DeClerck2024], while the present work provides the first systematic investigation of $\mathrm{MonoDC}$ and a direct comparison between alternative leader-based constructions.
+Moreover, the proposed approach does not rely on platform-specific features, but rather on general interaction patterns between content producers and their audiences, making its application to other platforms–where visibility and influence may be structured through different mechanisms–a natural direction for future research.
+Competing interests
+The authors declare that they have no competing interests.
+Author’s contributions
+S.G.
+and F.S.
+conceived and designed the study.
+F.S.
+and A.M.
+collected the data.
+S.G., F.S.
+and A.M.
+performed the modeling and data analysis with input from G.C.
+S.G and A.M.
+contributed to data preprocessing and visualization.
+S.G.
+and F.S.
+interpreted the results and drafted the manuscript.
+All authors contributed to revisions and approved the final version of the manuscript.
+Acknowledgements
+The authors are thankful to Lorenzo Federico for interesting discussions.
+Funding
+S.G.
+and F.S.
+were partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 “Education and Research” - Component C2 - Investment 1.1 - Next Generation EU “Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9.
+Appendix A Maximum Entropy Null Models
+We start from a real network $G^{*}$ (in the following, all quantities related to the observed network will be denoted by an asterisk *).
 Firstly, let us define a set $\mathcal{G}$ – the ensemble– including all possible graphs having the same number of nodes $N$ as in $G^{*}$ [^1] Since the number of nodes is fixed in the randomization, we drop the asterisk * from $N$..
@@
 $\langle\vec{C}\rangle=\vec{C}(G^{*}).$ (2)
-It can be shown that the condition in Eq. (2) can also be obtained through the maximization of the (log-)likelihood [38].
+It can be shown that the condition in Eq. (2) can also be obtained through the maximization of the (log-)likelihood [Garlaschelli2008].
 The framework described above is particularly flexible and can be adapted to different kinds of networks.
-In the present paper, we are interested in its bipartite versions for bipartite undirected and bipartite directed networks, i.e.
-respectively, the Bipartite Configuration Model (or BiCM [40]) and the Bipartite Directed Configuration Model (or BiDCM [41, 18]).
+In the following, we are interested in its bipartite versions for bipartite undirected and bipartite directed networks, i.e.
+respectively, the Bipartite Configuration Model (or BiCM [Saracco2015]) and the Bipartite Directed Configuration Model (or BiDCM [vanLidth2019, Becatti2019d]).
 In both cases, the constraints imposed in the entropy maximisation are, respectively, the undirected and directed degree sequences.
-More details about BiCM and BiDCM can be found in Appendices A.1 and A.2.
-2.2 $\mathrm{MonoDC}$ and $\mathrm{BiDC}$
-In the following, the two methods that we are proposing are presented.
-They both leverage the presence of leading content creators, i.e.
-users mostly posting original messages and less inclined to repost the messages created by others.
-There are multiple ways to define leading content creators: at the present level, we prefer to remain as general as possible.
-Therefore, in the following those users are called $\top$-⊤ -users, while the $\top$ users are called $\bot$-⊥ -users.
-2.2.1 $\mathrm{MonoDC}_{\top}$
-At least in the Italian political debate, it is common to observe that less relevant political figures are particularly active in liking and reposting the content created by more central leaders [^2] https://www.ilpost.it/2024/10/31/post-giorgia-meloni-social-commenti-parlamentari-fdi/.
-There could be multiple reasons for such a behaviour: intercept the audience of the leaders, demonstrate loyalty to either the leaders, the party or the objective of the campaign, or simply to contribute to the campaign of the leaders by increasing the visibility of their content.
-Therefore, it is reasonable to expect the activity of politicians to be much more focused on the activity of their party than other users.
-To capture such a phenomenon, the MonoDC studies the interactions among content creators through their activity in creating new contents or reposting the one created by others.
-The steps of MonoDC are represented as a cartoon in Fig.
-1.
-We describe data from Twitter/X as a bipartite directed network between users and posts in which edges from users to posts indicate authorships, while the vice versa indicates retweets.
-The information contained in the directed bipartite network is then projected to describe the retweet network, i.e.
-a directed multilink network in which nodes are users and the number of edges from user $A$ to user $B$ is the number of retweets that $B$ made to the post written by $A$.
-In this representation, we expect the (bipartite directed) degrees to play a role, since they represent the activity of users in either creating new posts or retweeting existing ones (represented, respectively, as the out- and in-degree in the bipartite directed network of the user nodes) and the virality of the various posts (i.e.
-the out-degree of the post nodes).
-Therefore, we can use a maximum entropy null model for directed bipartite networks as a statistical benchmark to validate the empirical retweet network, i.e.
-the Bipartite Directed Configuration Model (BiDCM) mentioned in Section 2.1.
-Technical details about BiDCM can be found in Appendix A.2.
-The result of such validation is a directed binary network, in which an edge represents that the retweet activity between the two nodes is non-trivial, i.e.
-it cannot be explained by just the attitude of users towards posting and reposting and the virality of the messages shared.
-We further restrict the validated network to only users in the set of content creators, i.e.
-the $\top$ user, in our jargon.
-The main idea is that the null model is not informed about the orientation of the users and therefore expects the debate to be “flat”, i.e.
-without any particular focus of the users towards the content created by some of the accounts.
-In this sense, communities should emerge naturally by comparing empirical data with the null model.
-Such a structure can be further captured by running a community detection algorithm on the validated retweet network.
-Finally, the labels obtained by the community detection are extended to $\bot$ users (i.e.
-the users that are not in the $\top$ set) by implementing Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network restricted to $\top$-users.
-2.2.2 $\mathrm{BiDC}_{\top}$
-The intuition behind $\mathrm{BiDC}_{\top}$ is to capture discursive communities through the perception of “standard” users of the similarity in the content created by the leaders of the debates [18, 43, 27, 19].
-Such a phenomenon is more stressed during political debates, being more or less intense depending on the electoral system of the country [43].
-Following the procedure outlined in Fig.
-2, we built the bipartite network between leading content creators and “standard” users: a link is present if the standard user has retweeted at least once the corresponding content creator.
-We then projected the bipartite network to a monopartite co-occurrences network among content creators.
-As in the case of $\mathrm{MonoDC}_{\top}$, we validate the network by comparing the empirical data with a maximum entropy null model for bipartite networks, i.e the Bipartite Configuration Model (BiCM) mentioned in Section 2.1.
-In this case, the information about the degree of the users is capturing the number of different retweeters and the number of different content creators retweeted for $\top$-⊤ - and $\bot$-⊥ -users, respectively; more details about the implemented maximum entropy null model can be found in Appendix A.1.
-BiCM, while accounting for the degree sequences of both layers, is not informed about the presence of communities and therefore expects the co-occurrence network not to have a modular structure.
-In this sense, any non-trivial focus of the users towards groups of content creators should be captured by comparing empirical data against the null model.
-Such a structure can be further refined by running a community detection algorithm on the validated projection network.
-Similarly to $\mathrm{MonoDC}_{\top}$, labels are extended to $\bot$ users by using Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network.
-2.3 Measuring the similarity of two partitions
-To assess the similarity between two ways of partitioning the same set of users, we use of the V-measure ($\mathrm{VM}_{\beta}$), a parametric measure of cluster similarity proposed by [31].
-Let $S(P)$ be the Shannon entropy of $P$, and $S(P\mid Q)$ be the conditional entropy of $P$ given $Q$. If $P$ and $Q$ are two partitions, the $\mathrm{VM}_{\beta}(P,Q)$ is defined as:
-$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)I(P,Q)}{S(P)+\beta S(Q)}$
-where $I(P,Q)$ is the mutual information between $P$ and $Q$.
-For $\beta=1$ we have
-$\mathrm{VM}_{1}(P,Q)=\frac{2I(P,Q)}{S(P)+S(Q)}=1-\frac{VI(P,Q)}{S(P)+S(Q)}$
-i.e., the $\mathrm{VM}_{1}$ accounts for the mutual information between the two partitions in a way that does not depend on their entropy, being equivalent to the Variation of Information [44], normalized by the total entropy of the partitions.
-With respect to other information-theoretic metrics, the $\mathrm{VM}_{1}$ and the VI have the advantage of being two distances – in particular, of satisfying the triangle inequality.
-Further, the $\mathrm{VM}_{1}$ always takes values in $[0,1]$, making it possible to directly compare the scores obtained for different combinations of algorithms and/or datasets.
-By letting $\beta$ vary in $[0,+\infty)$, we consider both scenarios in which one partition consists of sub-communities (e.g., political movements/trends within parties) or super-communities (e.g., political coalitions) of the other one.
-More details about the V-measure can be found in Appendix B.
-3 Results
-3.1 Data and networks
-To compare our approach for discursive community with standard community detection algorithms, we collected Twitter data regarding the three main political events debated on Italian Twitter in 2022.
-For each such topic of discussion, we identified a set of relevant keywords that were used to collect tweets using Twitter’s Academic API.
-Let us remark that data were downloaded before user’s verification became a feature of the X’s Premium subscription [^3] https://en.wikipedia.org/wiki/Twitter_verification..
-A summary of the datasets considered in this paper is reported in Table 1.
-Table 1: A summary of the datasets considered in the present paper
-Dataset Unique users Unique tweets Annotated users Verified users Influential users H-$\geq$3 users
-President1 119.018 711.497 292 1148 5804 1296
-Crisis2 113.876 1.414.181 229 1004 7038 2554
-Elections3 222.356 2.784.951 239 1555 12184 4382
-  • 1
-    Debate around the election of the President of the Republic. The political debate ended with the re-election of Sergio Mattarella on January 29, 2022.
-  • 2
-    Debate around the political crisis that led to the fall of the Draghi government on July 21, 2022.
-  • 3
-    Debate around the Italian general elections held on September 25, 2022, that led to Giorgia Meloni becoming the new Prime Minister.
-3.2 The benchmark
-In general, there is no reason to expect data to cluster due to some metadata [45].
-Multiple effects can cause –especially in online social networks– different users to interact with the same content creators.
-Nevertheless, due to the strictly political nature of the debates captured by our data, we expect to observe a signal of the political orientation of the users in the cluster that can be used as a benchmark for the classification methods under study.
-To construct our benchmark labeling, we first focused on just the verified accounts of Italian politicians and political parties.
-We choose verified accounts since their identities are certified by the platform.
-We manually labeled these accounts according to their political party and then grouped these labels according to the coalition each party belonged to during the general elections.
-This procedure yielded two annotation levels, which we refer to as “parties” and “coalitions”.
-To check if this benchmark is well-posed, we studied the retweet network restricted to the annotated users.
-Retweets are usually considered as a sort of endorsement of the content shared [1], so we expected users with a similar political orientation to cluster in this network.
-The modularity of the partitions induced by the annotations, whose values are all above 0.6 (cfr.
-Table 2), confirms our intuition and supports the idea of measuring the quality of a community detection algorithm also based on its ability to retrieve our manual annotations.
-Table 2: Retweet networks of annotated users and modularity (Q) of the partition induced by the manual annotation at the party and coalition level.
-Dataset Nodes Links $Q(\text{parties})$ $Q(\text{coalitions})$
-President 169 219 0.763 0.618
-Crisis 145 246 0.721 0.641
-Elections 154 355 0.777 0.649
-To compare different algorithms not just based on the small minority of users that we manually annotated, we extended the annotations to the rest of the network using a well-known Label propagation algorithm [42] –taking the most frequent label out of 100 independent realizations.
-These propagated labels cannot be considered equally accurate, but they arguably represent the best possible guess of a user’s political leaning, which can be made only based on how that user is connected to verified politicians in the retweet network.
-Let us notice that such an approach favors our models, in principle, since we are using the same label propagation algorithm to label standard users.
-In a sense, if we correctly tag $\top$ users, we will implicitly also meet the benchmark for $\bot$ users.
-It is indeed true, but it is an honest way to propagate the labels of verified accounts, as we cannot trust, in general, the one of other accounts that could be present in the user description.
-3.3 Limits of standard methods
-Let us start by considering the partitions obtained running four off-the-shelf community-detection algorithms on the entire retweet network: (i) Louvain, a heuristic for modularity maximization; (ii) Label propagation, a message passing algorithm where each node tries to align with the majority of its neighbors; (iii) Infomap, which is based on finding the optimal encoding for random walks on the graph; (iv) Inference, whose rationale is inferring the parameters of a stochastic block-model through a maximum-likelihood approach.
-For Louvain, Label propagation and Infomap, we used the implementation available in igraph (igraph.org, Python interface).
-For Inference, we used the minimize_blockmodel_dl() function provided in graph-tool (graph-tool.skewed.de), that fits a stochastic block-model to the given graph by minimizing the model’s description length.
-To guide the inference process, we set it explicitly to search for a planted-partition, that is, exclusively for assortative communities.
-In addition, we refined the algorithm as recommended in graph-tool’s documentation, that is, by running 100 times the multiflip_mcmc_sweep() method with zero temperature and 100 iterations.
-We ran each algorithm 100 times [^4] Each single run of the Louvain algorithm is already obtained taking the partition with the greatest modularity out of 100 independent executions with shuffled vertices.
-Similarly, each run of the Label propagation algorithm is already obtained by taking, for each node, the most frequent label out of 100 independent executions.
-and measured: the $\mathrm{VM}_{1}$ between each partition, restricted to the annotated users, and the annotations, both at the party and coalition level; the pairwise $\mathrm{VM}_{1}$ between partitions obtained with different algorithms, restricted to the verified users.
-Fig.
-3 reports these $\mathrm{VM}_{1}$ values, considering only the parties annotations.
-We focused on party annotations, but similar results are observed with other algorithms and with coalitions.
-Fig.
-3 highlights the limited accuracy in recognizing the affiliation of renowned politicians –$\mathrm{VM}_{1}\in[0.4,0.6]$– of the communities assigned to verified users by different off-the-shelf algorithms.
-Furthermore, as shown in Fig.
-10 in Appendix C, standard algorithms show great variability of the communities assigned to verified users by different off-the-shelf algorithms –pairwise $\mathrm{VM}_{1}\in[0.3,0.7]$.
-Overall, these findings motivate the research for a different approach to the detection of discursive communities on Twitter/X.
-Figure 3: Limits of standard methods.
-The $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms, restricted to the annotated users, and the parties’ annotations.
-Standard methods show a limited accuracy in recognising the affiliation of renowned politicians.
-3.4 $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$
-We now show the results of our methods, i.e.
-$\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, when we choose verified users as the class of content creators: in this case, we will write $\top=V$.
-In the following, the communities on the network of $V$ users are obtained by running the Louvain algorithm on the projected network and are then propagated to the rest of the network.
-We tested our community detection algorithm, as the Label propagation and Infomap: their results are very similar to those obtained with Louvain ($\mathrm{VM}_{1}>0.8$) and yield similar results also for both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$.
-Communities obtained with the inference-based method are instead way less accurate in reconstructing the annotations.
-3.4.1 Comparison with standard methods
-In Fig.
-4, we use the $\mathrm{VM}_{1}$ to show how accurately each method reconstructs the manual annotations at the party and coalition level, respectively in the first and second row.
-Each violin represents the distribution over 100 independent runs, and we compare $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with the partitions obtained running the Louvain algorithm on the entire retweet network –which performs no worse than other algorithms (cfr.
-Fig.
-3).
-$\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions.
-The fact that in all three cases the community detection algorithm used is Louvain signifies that what ultimately matters is focusing on the right network.
-For completeness, the comparison between the performances of our methods and Louvain is presented in Fig.
-11 in Appendix C.
-Even in this case, both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ outperform Louvain.
-Nonetheless, let us note that, when comparing the results of the propagated labels of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}$ with the annotations of verified users propagated using the same label propagation algorithm, our methods have an advantage.
-Nevertheless, the performances of standard methods are worse, even when limited to annotated users.
-Figure 4: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods.
-The $\mathrm{VM}_{1}$ between partitions of the annotated users, obtained with different methods, and the annotations at party (first row) and coalition (second row) level.
-$\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions.
-It could be argued that standard algorithms, when executed on the entire retweet network, return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect.
-Fig.
-5 shows that this is not the case: when compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e.
-at essentially all scales at which the political spectrum can be observed.
-Figure 5: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods at different resolution levels.
-The $\mathrm{VM}_{\beta}$ between partitions of the annotated users, obtained with different methods, and the annotations, for different values of $\beta$ and both at the parties (top) and coalitions (bottom) level.
-Each point is the average over 100 runs of the method indicated by the colour and marker.
-Analogous results are obtained when considering the propagated annotations.
-When compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e.
-at essentially all scales at which the political spectrum can be observed.
-In this sense, standard algorithms, when executed on the entire retweet network, do not return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect.
-In Fig.
-6, the Sankey diagrams corresponding to each evaluated method provide a detailed depiction of the political composition of the extracted communities.
-$\mathrm{MonoDC}_{V}$ exhibits remarkable precision in delineating party structures —successfully distinguishing individual parties such as Azione and Italia Viva, which presented a unified list during the elections— and further reveals latent political currents within party boundaries.
-Conversely, $\mathrm{BiDC}_{V}$ proves highly effective in reconstructing the official electoral coalitions, achieving a near-exact correspondence with their institutional configurations.
-The Louvain algorithm, applied to the entire retweet network, yields a substantially noisier community structure that conflates partisan and coalition-level affiliations.
-Notably, Forza Italia emerges as an isolated community, distinct from its allied parties, while several political figures are clustered within communities dominated by their ideological adversaries.
-(a) Louvain on RT network.
-(b) $\mathrm{MonoDC}_{V}$.
-(c) $\mathrm{BiDC}_{V}$.
-Figure 6: Qualitative comparison of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with Louvain on the Elections dataset.
-For each of the considered methods, the Sankey diagrams show the composition of the extracted communities in terms of political parties.
-$\mathrm{MonoDC}_{V}$ very accurately detects parties (Azione and Italia Viva ran together as a single list at the elections), plus some political movements/trends within parties; $\mathrm{BiDC}_{V}$ almost perfectly detects the official political coalitions of those elections; Louvain, executed on the entire retweet network, provides a more noisy classification that is a mix of the two levels, with Forza Italia separated from its coalitions and many politicians contributing to communities led by some of their rivals.
-Finally, since $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ capture the political leaning of users at different scales, in Fig.
-7 we use the $\mathrm{VM}_{\beta}$ to check whether the clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$.
-This seems to be true, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
-Figure 7: Different resolution of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$.
-The $\mathrm{VM}_{\beta}$, for variable $\beta$, between partitions of all users obtained running $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$.
-The clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
-3.4.2 Final remarks
-Both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ benefit from a pre-processing stage capable of extracting, from the original retweet data, a more informative network of just such leading content creators.
-The role that statistical validation plays in the two cases is, however, incomparable.
-In $\mathrm{MonoDC}_{V}$, the validated retweet networks of verified accounts of politicians include 94-97% –depending on the dataset– of all edges present before the validation.
-The validation is performed with respect to the patterns emerged from the analysis of the entire network and the fact that almost all retweets made by politicians are statistically significant means that their activity on X is extremely focused on just the content produced by a limited niche of users.
-Since these users are their party’s leaders and fellow members, $\mathrm{MonoDC}_{V}$ is the approach that best captures the loyalty to the party and, hence, the difference between parties that apparently occupy a similar position in the political spectrum.
-The price to pay for this finer classification is its minor stability, as can be seen in Fig.
-4.
-In this sense, $\mathrm{MonoDC}_{V}$ seems particularly promising to examine the mesoscale structure of such a subgraph to infer the main poles of the political discussion, since the phenomenon we are leveraging on is particularly strong.
-Differently from the $\mathrm{MonoDC}_{V}$’s case, for $\mathrm{BiDC}_{V}$ the impact of the statistical validation is critical: the number of links in the validated network is 7-12% –depending on the dataset– of those present in the non-validated projection.
-When the political sides are clear, this validated signal coming from the audience returns a very stable partition (cfr.
-Fig.
-4).
-In addition, $\mathrm{BiDC}_{\top}$ has the advantage that the similarity between any two content creators $u,v\in\top$ can be established even when $u$ and $v$ did not directly interact, making $\mathrm{BiDC}_{\top}$ more flexible, robust and widely usable than $\mathrm{MonoDC}_{\top}$.
-For the sake of completeness, we ran the same set of community-detection algorithms considered in Section 3.3 on the non-validated network, as well as on the validated projections with different significance levels [^5] To properly target the significance level in validating the projected network, we implemented the False Discovery Rate (FDR) procedure [46], see Appendices A.1.1 and A.2.1., cfr.
-Fig.
-12 in Appendix C.
-The results on the non-validated network are very poor, whereas the significance level does not seem to have a great impact on the inferred similarity network.
-3.5 Other classes of content creators
-In the previous subsections, we examined the behaviours of the $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ after identifying verified users as relevant content creators, i.e.
-the main leaders of the shared discourses.
-Verified users ($V$) are a very natural choice for the $\top$ layer of both $\mathrm{MonoDC}_{\top}$ –because their behavior on the platform is bound to their role outside of it– and $\mathrm{BiDC}_{\top}$ –because they are generally known to and easily categorizable by the audience.
-There are other ways, however, to select users based on their actual or presumed influence on online discussions.
-For instance, [4] defined “influential” users those who have more followers than friends (i.e.
-the accounts a user follows, sometimes referred to as “followees” in the literature), and who, in the debate under study, received more mentions than those they made.
-Another way to measure the impact of a user in a community, commonly employed for scientists and academics, is the Hirsch index [47] –shortly, H-index– which can be adapted to X’s mode of operation by saying that a user $u$’s H-index in the debate is the largest $h$ such that at least $h$ tweets authored by $u$ received at least $h$ retweets each.
-In the following, we use $I$ to denote the set of influential users and $H$ for the set of top H-index users (defined as users having H-index3absent3$\geq$ 3≥ 3, see Fig.
-13 in Appendix C for a justification of the threshold).
-We aim to assess both what happens when $\top=I$ or $\top=H$ is used in $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, and how stable and accurate is the classification obtained for these two sets of central users when compared with the entire dataset.
-(a) $\mathrm{MonoDC}_{\top}$, for different choices of $\top$, compared with annotations at the parties level.
-(b) $\mathrm{BiDC}_{\top}$, for different choices of $\top$, compared with annotations at the coalitions level.
-Figure 8: Comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$.
-The $\mathrm{VM}_{1}$ between partitions obtained with different choices of $\top$ and the annotations, for the Elections dataset.
-For each method, we used the level of annotation that the method more accurately reproduces, that is, parties for $\mathrm{MonoDC}_{\top}$ (a) and coalitions for $\mathrm{BiDC}_{\top}$ (b).
-The columns refer to the user set upon which the VM is computed.
-For instance, “user set = influential” means that in the second column we only consider how influential users have been labeled by different methods, temporarily ignoring the rest of the network.
-In Fig.
-8, the $\mathrm{VM}_{1}$ is used to compare partitions obtained with different choices of $\top$ with the (propagated) annotations ($x$-axis).
-Here we consider the Elections dataset, but similar results hold for the other datasets.
-In Appendix C), we also compare the cases $\top=H$ and $\top=I$ with the case $\top=V$, taken as a reference.
-The choice of the $\top$ set has an impact on both $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$.
-Especially for $\mathrm{BiDC}_{\top}$, $\top=V$ yields the best overlap with the annotations.
-It could be argued that it is due to $V$ being a small set that includes all manually annotated users, but $H$ outperforms $I$ despite the former includes just 20-50% (depending on the dataset) of the annotated public figures, compared with 80-90% of the latter.
-This supports the claim that $V$ users are the key to understanding the sides of the public discourse.
-4 Conclusions
-In this paper, we focused on X/Twitter and on so-called discursive communities, i.e.
-the sub-groups that naturally emerge in Online Social Networks (OSNs) as users cluster around shared narratives, language, or value systems during public debates.
-What we observe in OSNs is a network of interactions, where edges exist for a combination of factors – e.g.
-the visibility, the status and the behaviour of different users on and off the platform, and the platform’s recommendation system – and where understanding what led to the creation of a community is almost as important as detecting the community itself.
-In this paper, we investigate the role of verified users in shaping discursive communities in the pre-Premium accounts era.
-Introduced in 2009, verified accounts were originally intended to prevent impersonation of public figures or organizations [21, 22].
-At first, the platform itself selected which accounts to verify, based on their relevance in public discourse.
-These accounts were marked by a blue checkmark next to their name.
-While verification helped authenticate the presence of public figures, it also introduced a hierarchy among users—one that arguably amplified the influence of verified accounts in online debates [48].
-Despite these concerns, in November 2022, following Elon Musk’s acquisition of Twitter, verification became a paid feature available through X’s Premium subscription [24].
-We conduct our analysis around three key political events in Italy during 2022: the election of the President of the Republic, the political crisis, and the subsequent national elections.
-Our initial observation is that a strong signal of political alignment is detectable in the data: when clustering verified politicians in the retweet network by political party, the resulting partition shows high modularity.
-We then explore the limitations of existing community detection algorithms.
-When comparing their output to manually annotated data, both descriptive and inferential methods often produce noisy and imprecise community structures.
-To address this, we evaluate the performance of two algorithms designed to identify discursive communities: $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, both of which begin by labeling verified users, given their central role in public debate.
-The $\mathrm{MonoDC}_{V}$ approach is grounded in a typical political dynamic, where less prominent figures tend to retweet content from party leaders.
-Conversely, the $\mathrm{BiDC}_{V}$ method—originally proposed by [18]—builds an undirected similarity network between content creators by analyzing audience overlap.
-In both cases, we reduce the effect of random noise—an inherent feature of OSNs [28, 29]—by comparing empirical data to a maximum entropy null model, which accounts for user activity [30].
-Both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ successfully reproduce the organization of political debates into discursive communities shaped by parties or alliances.
-Notably, using verified users as leading content creators consistently outperforms alternative selections based solely on user activity.
-This suggests that the status conferred by verification plays a more decisive role in structuring the online conversation than activity alone.
-In summary, our work offers two take-home messages.
-First, in X/Twitter datasets from before November 2022, leveraging verified users allows for a clearer and more accurate identification of discursive communities.
-Alternative approaches relying solely on user activity result in noisier partitions, suggesting that user status is a major factor in shaping the online debate.
-Second, our findings are based on comparisons conducted before account verification became a paid feature.
-This raises concerns about the potential consequences of monetized verification: by enabling users to purchase visibility and prestige, the platform may open the door to malicious actors seeking to manipulate discourse and distort the public agenda.
-A deeper comparison between “old” –i.e.
-pre-November 2022– and “new” –i.e.
-post-November 2022– verified users will be the target of further research.
-Competing interests
-The authors declare that they have no competing interests.
-Acknowledgements
-The authors are thankful to Lorenzo Federico for interesting discussions.
-SG and FS were partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 “Education and Research” - Component C2 - Investment 1.1 - Next Generation EU “Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9.
-Appendix A Maximum entropy null models for bipartite networks
-A.1 Bipartite Configuration Model
-The framework described in Section 2.1 can be adapted to bipartite networks.
-In particular, in the following we will use the degree sequences of nodes of both layers as constraints: in this case, the null-model takes the name of Bipartite Configuration Model (or BiCM [40]).
-Let us call the two layers of a bipartite network $\top$ and $\bot$ and let $N_{\top}$ and $N_{\bot}$ be their dimensions; nodes belonging to the $\top$-⊤ -layer will be denoted by Latin indices, while the ones belonging to the $\bot$-⊥ -layer will be denoted by Greek ones.
+A.1 Bipartite Configuration Model (BiCM)
+Given a bipartite network with layers $\top$ (e.g., users) and $\perp$ (e.g., tweets), let $N_{\top}$ and $N_{\bot}$ be their dimensions.
+Nodes belonging to the $\top$-layer will be denoted by Latin indices, while the ones belonging to the $\bot$-layer will be denoted by Greek ones.
 Any bipartite network is completely described by its biadjacency matrix $\mathbf{B}$, i.e.
@@
 the number of connections that the nodes have, is simply $k_{i}=\sum_{\alpha}b_{i\alpha}$ ($h_{\alpha}=\sum_{i}b_{i\alpha}$).
-Following the procedure sketched in Section 2.1 returns that the probability of observing the generic (bipartite) graph $G_{\text{Bi}}$ is
-$\begin{split}P(G_{\text{Bi}})=&\prod_{i,\alpha}\dfrac{e^{-(\theta_{i}+\eta_{\alpha})b_{i\alpha}(G_{\text{Bi}})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}\end{split}$ (3)
+The Bipartite Configuration Model preserves the degree sequences $k_{i}=\sum_{\alpha}b_{i\alpha}$ and $h_{\alpha}=\sum_{i}b_{i\alpha}$.
+Following the procedure sketched above, the probability of observing a graph $G_{\text{Bi}}$ in the ensemble is:
+$P(G_{\text{Bi}})=\prod_{i,\alpha}\frac{e^{-(\theta_{i}+\eta_{\alpha})b_{i\alpha}(G_{\text{Bi}})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}},$
 where $\vec{\theta}$ and $\vec{\eta}$ are the Lagrangian multipliers associated with the degree sequences on, respectively, layers $\top$ and $\bot$.
-Eq.
-(3) indicates that probabilities per graph factorize in terms of independent probabilities per link $p_{i\alpha}=\frac{e^{-(\theta_{i}+\eta_{\alpha})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}$.
-To get the numerical values of the Lagrangian multipliers, we can impose the likelihood maximization [38].
-In the present analysis, we used the python module bicm to calculate the numerical values of the probabilities: bicm, described in Ref. [49] as a submodule of NEMtropy, is also a standalone module.
-A.1.1 BiCM validation of the monopartite projection of a bipartite network
-Using BiCM as a benchmark permits to validate the co-occurrences in the real network [50].
+The corresponding link probabilities are:
+$p_{i\alpha}=\frac{e^{-(\theta_{i}+\eta_{\alpha})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}.$
+To get the numerical values of the Lagrangian multipliers, we can impose the likelihood maximization [Garlaschelli2008].
+A.1.1 Validated Projection under BiCM
+Using BiCM as a benchmark permits to validate the co-occurrences in the real network [Saracco2017].
 In fact, observed co-occurrences are simply the “projection” of the information contained in the bipartite network into one of the two layers, say $\top$:
@@
 (in Ref.
-[40] co-occurrences were also called V-motifs since, if the two layers of the bipartite networks are drawn horizontally, they sketch a “V” between the layers).
+[Saracco2015] co-occurrences were also called V-motifs since, if the two layers of the bipartite networks are drawn horizontally, they sketch a “V” between the layers).
 Since in BiCM the probabilities per link are independent, then
@@
 Calculating exactly the Poisson-Binomial distribution is computationally costly.
-It was shown, that when probabilities are relatively small –as in our case–, the Poisson-Binomial distribution can be finely approximated with a Poisson distribution [51].
+It was shown, that when probabilities are relatively small –as in our case–, the Poisson-Binomial distribution can be finely approximated with a Poisson distribution [Hong2013].
 Therefore, it is possible to efficently calculate the p-value associated with each observed co-occurrence $(V^{ij})^{*}$.
 Since the number of different co-occurrences is $\binom{N_{\top}}{2}$, we need multiple hypothesis testing.
-In the present manuscript, we implemented the False Discovery Rate (FDR,  [46]) since it controls the rate of False Positives.
+In the present manuscript, we implemented the False Discovery Rate (FDR,  [Benjamini1995]) since it controls the rate of False Positives.
 Again, also the validated projection used in the present manuscript was calculated using the python module bicm.
-A.2 Bipartite Directed Configuration Model
+A.2 Bipartite Directed Configuration Model (BiDCM)
 Let us now consider the case in which we have a bipartite directed network and we want our model to include the information about the in- and out-degrees of nodes on both layers.
-It can be shown that the probability per bipartite directed graph $G_{\text{BiD}}$ factorises in terms of the BiCM probability per two undirected bipartite graphs $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, describing, respectively, all edges from layer $\top$ to layer $\bot$, and vice versa [41, 18]:
+It can be shown that the probability per bipartite directed graph $G_{\text{BiD}}$ factorises in terms of the BiCM probability per two undirected bipartite graphs $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, describing, respectively, all edges from layer $\top$ to layer $\bot$, and vice versa [vanLidth2019, Becatti2019d]:
 $P_{\text{BiDCM}}(G_{\text{BiD}})=P_{\text{BiCM}}(G_{\text{Bi}}^{\top\rightarrow\bot})\cdot P_{\text{BiCM}}(G_{\text{Bi}}^{\top\leftarrow\bot}).$
@@
 $\begin{split}q_{i\alpha}^{\top\rightarrow\bot}=&\dfrac{e^{-(\theta_{i}^{\text{out}}+\eta_{\alpha}^{\text{in}})}}{1+e^{-(\theta_{i}^{\text{out}}+\eta_{\alpha}^{\text{in}})}};\\ q_{i\alpha}^{\top\leftarrow\bot}=&\dfrac{e^{-(\theta_{i}^{\text{in}}+\eta_{\alpha}^{\text{out}})}}{1+e^{-(\theta_{i}^{\text{in}}+\eta_{\alpha}^{\text{out}})}}.\end{split}$
-The numerical values of the $\theta_{i}^{\text{out}}$, $\theta_{i}^{\text{in}}$, $\eta_{\alpha}^{\text{out}}$ and $\eta_{\alpha}^{\text{in}}$ can be obtained through the maximization of the likelihood [38].
-A.2.1 BiDCM validation of the monopartite projection of a bipartite network
-Following the procedure described in Appendix A.1.1, we can provide a statistical validation of the monopartite multilink directed network obtained by projecting, say, on layer $\top$, a bipartite directed network.
+The numerical values of the $\theta_{i}^{\text{out}}$, $\theta_{i}^{\text{in}}$, $\eta_{\alpha}^{\text{out}}$ and $\eta_{\alpha}^{\text{in}}$ can be obtained through the maximization of the likelihood [Garlaschelli2008].
+A.2.1 Validated Projection under BiDCM
+As done for the BiCM, we can provide a statistical validation of the monopartite multilink directed network obtained by projecting, say, on layer $\top$, a bipartite directed network.
 The number of edges from node $i$ to node $j$ in the projection is
@@
 $\langle\mathcal{V}^{ij}\rangle=\sum_{\alpha}q_{i\alpha}^{\top\rightarrow\bot}q_{j\alpha}^{\top\leftarrow\bot}.$ (3)
-A.2.2 The application of BiDCM-validated projection to the bipartite network of users and posts
+A.2.2 BiDCM-Validated Projection for Users and Posts
 In the analyses above, the BiDCM-validated projection was used to validate the multilink retweet network among users.
@@
 $q_{i\alpha}^{\top\rightarrow\bot}=\dfrac{e^{-(\theta_{i}^{\text{out}}+\eta^{\text{in}})}}{1+e^{-(\theta_{i}^{\text{out}}+\eta^{\text{in}})}};$
-in this case, the Bipartite Configuration Model reduces to a Bipartite Partial Configuration Model [50].
+in this case, the Bipartite Configuration Model reduces to a Bipartite Partial Configuration Model [Saracco2017].
 Imposing that the constraints, i.e.
@@
 (5) is remarkable: the expected value $\langle\mathcal{V}^{ij}\rangle$ can be easily calculated without knowing explicitly the value of the Lagrangian multipliers.
-Therefore, p-values against BiDCM of the observed $\mathcal{V}^{ij}$ can be easily calculated, and further statistically validated using the FDR procedure described above [46].
-A.2.3 Limits of applicability
+Therefore, p-values against BiDCM of the observed $\mathcal{V}^{ij}$ can be easily calculated, and further statistically validated using the FDR procedure described above [Benjamini1995].
+A.2.3 Limitations
 Besides the simplicity of Eq.
-(6), there is an important issue, already mentioned for the present application in Ref.
-[18].
+(5), there is an important issue, already mentioned for the present application in [Becatti2019d].
 In the null model implemented above, constraints are satisfied on average.
@@
 Such an issue is going to be the target of further research.
-Appendix B The V-measure for clustering similarity
-Let us define, respectively the homogeneity and the completeness of $Q$ with respect to $P$ as
-$h(P,Q)$ $=1-\frac{S(P\mid Q)}{S(P)}$
-$c(P,Q)$ $=1-\frac{S(Q\mid P)}{S(Q)}.$
-Note that $h=1$ if and only if the clusters obtained with $Q$ are sub-clusters of those obtained with $P$, whereas $c=1$ if and only if the clusters obtained with $Q$ are super-clusters of those obtained with $P$.
-It can be shown that the V-measure $\mathrm{VM}_{\beta}(P,Q)$ is the harmonic mean of $h$ and $c$, weighted by the parameter $\beta$ [31]:
-$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)h(P,Q)c(P,Q)}{\beta h(P,Q)+c(P,Q)}.$
-It is worth noting that the VM of two partitions does not decrease linearly with the number of elements on which the two partitions disagree: the $\mathrm{VM}_{1}$ between two partitions obtained with the same algorithm is generally not greater than 0.9; if we consider two partitions that differ by only a small percentage of the network that is randomly relabeled, their $\mathrm{VM}_{1}$ is about 0.8 with $\sim$5of the network relabeled, and decreases to about 0.6 when $\sim$10-13of the network has been relabeled (cfr.
-Fig.
-9 in Appendix B).
-Based on our tests, we deduce that a reasonably good $\mathrm{VM}_{1}$ between two partitions obtained with different methods/algorithms lies in the range $[0.6-0.9]$.
+A.3 Implementation
+In the paper, we used the implementation of the BiCM and BiDCM available in the python module NEMtropy, described in [Vallarano2021].
+bicm is also available as a standalone module.
+As a reference, the running time for BiDCM on the largest dataset are reported in Table 3.
+Table 3: Mean and standard deviation of the execution time, expressed in seconds, of the different steps of $\mathrm{BiDC}_{V}$ on the Elections dataset, the largest dataset considered in this paper.
+These timings highlight that the cost of $\mathrm{BiDC}_{V}$ is dominated by the label propagation algorithm presented in [Raghavan2007a].
+mean std
+algorithm step
+Label propagation on rt network 943.76 62.14
+Louvain on top network 2.64 0.03
+validated projection p=0.01 69.74 2.42
+Appendix B V-Measure for Clustering Similarity
+To assess the similarity between two ways of partitioning the same set of users, we use of the V-measure ($\mathrm{VM}_{\beta}$), a parametric measure of cluster similarity proposed in [rosenberg2007v].
+Let $S(P)$ be the Shannon entropy of $P$, and $S(P\mid Q)$ be the conditional entropy of $P$ given $Q$. If $P$ and $Q$ are two partitions, the $\mathrm{VM}_{\beta}(P,Q)$ is defined as:
+$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)I(P,Q)}{S(P)+\beta S(Q)}$
+where $I(P,Q)$ is the mutual information between $P$ and $Q$. $VM_{\beta}$ can be interpreted as the harmonic mean, weighted by $\beta$, of homogeneity $h$ and completeness $c$:
+$h(P,Q)=1-\frac{S(P|Q)}{S(P)},\quad c(P,Q)=1-\frac{S(Q|P)}{S(Q)},$
+$VM_{\beta}(P,Q)=\frac{(1+\beta)hc}{\beta h+c}.$
+Note that $h=1$ if and only if the clusters obtained with $Q$ are sub-clusters of those obtained with $P$, whereas $c=1$ if and only if the clusters obtained with $Q$ are super-clusters of those obtained with $P$ [rosenberg2007v].
+For $\beta=1$ we have
+$\mathrm{VM}_{1}(P,Q)=\frac{2I(P,Q)}{S(P)+S(Q)}=1-\frac{VI(P,Q)}{S(P)+S(Q)}$
+i.e., the $\mathrm{VM}_{1}$ accounts for the mutual information between the two partitions in a way that does not depend on their entropy, being equivalent to the Variation of Information [Meila2007], normalized by the total entropy of the partitions.
+With respect to other information-theoretic metrics, the $\mathrm{VM}_{1}$ and the VI have the advantage of being two distances – in particular, of satisfying the triangle inequality.
+Further, the $\mathrm{VM}_{1}$ always takes values in $[0,1]$, making it possible to directly compare the scores obtained for different combinations of algorithms and/or datasets.
+By letting $\beta$ vary in $0,+\infty)$, we consider both scenarios in which one partition consists of sub-communities (e.g., political movements/trends within parties) or super-communities (e.g., political coalitions) of the other one.
+It is worth noting that the VM of two partitions does not decrease linearly with the number of elements on which the two partitions disagree.
 Fig.
@@
 We consider two choices for the initial labeling: 5 equally sized random communities and the communities obtained running the Louvain algorithm.
-In both cases, the $\mathrm{VM}_{1}$ rapidly decreases to about 0.8 when $\sim$5of the network has been relabeled, and to 0.6 when $\sim$10-13of the network has been relabeled.
+In both cases, the $\mathrm{VM}_{1}$ rapidly decreases to about 0.8 when $\sim$5% of the network has been relabeled, and to 0.6 when $\sim$10-13% of the network has been relabeled.
 We also show that two different runs of the Louvain algorithm generate partitions whose $\mathrm{VM}_{1}$ is about 0.9, on average.
+Based on these tests, we deduce that a reasonably good $\mathrm{VM}_{1}$ between two partitions obtained with different methods/algorithms lies in the range [0.60.9][0.6-0.9].
 Figure 9: Understanding the values of $\mathrm{VM}_{1}$.
 The plot shows $\mathrm{VM}_{1}$ between independent runs of the Louvain algorithm compared with the $\mathrm{VM}_{1}$ between some partition of the Elections dataset –either 5 equally sized random communities, or the communities obtained running the Louvain algorithm– and a noisy version of the same partition obtained randomly reallocating part of the users, as a function of the percentage of the network that is reallocated.
-Appendix C Additional analyses
-In the following, we report additional figures and tables that have been left out of the main sections.
-For the sake of readability and clarity, all discussions can be found in the respective captions.
+Appendix C Data collection and political context
+Data were downloaded using Twitter/X official APIs, using a Researcher account.
+Data were collected using the Python module twarc2 (https://twarc-project.readthedocs.io/en/latest/twarc2_en_us/), using the search --archive command to access the archive of data.
+Table 4: Description of the queries used for the data collection
+Dataset Keywords Start time End time
+President quirinale 2022-01-01 2022-01-31
+presidente della repubblica
+mattarella
+Crisis crisidigoverno 2022-07-01 2022-08-01
+crisigoverno
+draghi
+governo
+Elections elezionipolitiche2022 2022-08-25 2022-10-02
+elezionipolitiche22
+elezioni
+PD
+Letta
+Calenda
+Azione
+Renzi
+ItaliaViva
+FdI
+Meloni
+Lega
+Salvini
+Conte
+M5S
+ForzaItalia
+Berlusconi
+In the President dataset, “Quirinale” is the name of the Roman hill where the headquarters of the Italian President of the Republic is.
+“Presidente della Repubblica” is the Italian name for President of the Republic, while “Mattarella” is the surname of the president that was confirmed in early 2022.
+In 2021, President Mattarella consistently declined to make himself available for a second term, despite sustained pressure from a broad range of political actors.
+Nonetheless, in the absence of an officially endorsed candidacy, he continued to attract a significant number of votes in successive ballots, obtaining 336 votes in the sixth round and 387 in the seventh.
+By 29 January, the failure of all alternative candidacies and the persistent coordination breakdown between the two major coalitions rendered the re-election of the incumbent a focal point of convergence.
+Following this deadlock, Mattarella accepted a renewed mandate and was re-elected with 759 votes.
+In Italian political discourse, “crisi di governo” is an idiomatic expression referring to a cabinet crisis, namely the fall of the executive without a systemic constitutional breakdown.
+Mario Draghi was the Prime Minister at the time of the crisis.
+On July 13, 2022, during a press conference, Conte, leader of the M5S, declared that his party would abstain from voting on the Aid Decree, which the government had designated as a matter of confidence.
+The following day, the Senate approved the decree, while M5S representatives withdrew from the chamber during the vote.
+Although this action did not constitute a formal withdrawal of governmental support, it was broadly interpreted as a substantial challenge to the administration’s agenda and effectively precipitated a political crisis within Draghi’s Cabinet.
+After strong debates in the two chambers of the Italian parliament, on July 21, Draghi definitively submitted his resignation to President Mattarella, who was compelled to accept, as the government’s loss of parliamentary support had by then become unequivocal and irreversible.
+The keywords used to gather tweets for the Elections dataset include the surname of the main political party leaders (i.e.
+Letta, Calenda, Renzi, Meloni, Salvini, Conte, Berlusconi), as well as the name of their party (PD, Azione, Italia Viva, Fratelli d’Italia, Lega, Movimento 5 Stelle, Forza Italia); other keywords relate to the event itself.
+Following the Draghi government’s collapse and resulting parliamentary deadlock, President Sergio Mattarella dissolved Parliament on July 21, 2022, calling early elections for September 25.
+The centre-right coalition led by Giorgia Meloni’s national-conservative Fratelli d’Italia (FdI) won an absolute parliamentary majority.
+Meloni was appointed Prime Minister on October 22, becoming Italy’s first female head of government.
+Appendix D Experimental Details and Extended Comparisons
+D.1 Community Detection Algorithms and Implementations
+Community detection algorithms used:
+  • Louvain, Label Propagation, Infomap from the igraph Python package.
+  • Stochastic Block Model (SBM) inference via graph-tool’s minimize_blockmodel_dl().
+The minimize_blockmodel_dl() function fits a stochastic block-model to the given graph by minimizing the model’s description length.
+To guide the inference process, we set it explicitly to search for a planted-partition, that is, exclusively for assortative communities.
+In addition, we refined the algorithm as recommended in graph-tool’s documentation, that is, by running 100 times the multiflip_mcmc_sweep() method with zero temperature and 100 iterations.
+We ran each algorithm 100 times.
+Each single run of the Louvain algorithm is already obtained taking the partition with the greatest modularity out of 100 independent executions with shuffled vertices.
+Similarly, each run of the Label propagation algorithm is already obtained by taking, for each node, the most frequent label out of 100 independent executions.
+D.2 Definition of $H$ Users
+We repeated MonoDC and BiDC using:
+  • Influential users ($I$): users with follower/followee ratio >1>1 and more mentions received than made.
+  • Top H-index users ($H$): users with $H\geq 3$ where $H$ is the largest number such that the user has at least $H$ tweets each with at least $H$ retweets.
+The threshold $H>3$ has been selected based on the elbow-method applied to the frequency of users for whom $H>h$ for different values of $h$ (cfr. Fig. 10).
+Figure 10: The distribution of the H-index in the three datasets.
+The threshold $H=3$ used to tell apart $\top$ and $\bot$ users is justified by the empirical H-index distribution having an elbow at 3; cumulatively, only about $1-2\%$ of the users pass the threshold.
+D.3 Effect of Significance Thresholds
+When running the $\mathrm{BiDC}_{V}$ method, we observed that only a small percentage of all edges resulted as being statistically significant.
+We therefore explored BiDCM validation at multiple significance levels ($\alpha=0.001$, $0.01$, $0.05$).
+Results, shown in Fig.
+11, were stable across values, confirming robustness.
+Non-validated projections performed poorly.
+Figure 11: An analysis of the role of validation ($\alpha$ denotes the significance level) in determining the partitions obtained for the $\top$ layer of $\mathrm{BiDC}_{V}$.
+Top: the $\mathrm{VM}_{1}$ between partitions of annotated $V$ users, obtained for different validation levels, and the annotations.
+Bottom: the $\mathrm{VM}_{1}$ between partitions of $V$ users, obtained for different validation levels, and those obtained with $\alpha=0.01$.
+Each point is a different run of the method indicated by the color and marker.
+The results on the non-validated network are very poor, whereas the significance level has a limited impact on the inferred partitions.
+D.4 Additional Comparisons
+As mentioned in the main article, we performed the following additional tests that overall confirm the main findings:
+  • A comparison of the partitions obtained with different off-the-shelf clustering algorithms, using the Louvain method as a benchmark (cfr. Fig. 12).
+  • A comparison of the partitions obtained with different methods for the entire network and the propagated annotations (cfr. Fig. 13).
+  • A cross-comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$ (cfr. Fig. 14).
 Figure 12: The average $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms and 100 runs of the Louvain algorithm (in green, the self-similarity of Louvain partitions).
@@
 Nonetheless, the performances of standard methods are worse, even when limited to annotated users.
-Table 3: Mean and standard deviation of the execution time, expressed in seconds, of the different steps of $\mathrm{BiDC}_{V}$ on the Elections dataset, the largest dataset considered in this paper.
-These timings highlight that the cost of $\mathrm{BiDC}_{V}$ is dominated by the label propagation algorithm presented in Ref.
-[42].
-mean std
-algorithm step
-Label propagation on rt network 943.76 62.14
-Louvain on top network 2.64 0.03
-validated projection p=0.01 69.74 2.42
-Figure 12: An analysis of the role of validation ($\alpha$ denotes the significance level) in determining the partitions obtained for the $\top$ layer of $\mathrm{BiDC}_{V}$.
-Top: the $\mathrm{VM}_{1}$ between partitions of annotated $V$ users, obtained for different validation levels, and the annotations.
-Bottom: the $\mathrm{VM}_{1}$ between partitions of $V$ users, obtained for different validation levels, and those obtained with $\alpha=0.01$.
-Each point is a different run of the method indicated by the color and marker.
-The results on the non-validated network are very poor, whereas the significance level has a limited impact on the inferred partitions.
-Figure 13: The distribution of the H-index in the three datasets.
-The threshold $H=3$ used to tell apart $\top$ and $\bot$ users is justified by the empirical H-index distribution having an elbow at 3; cumulatively, only about $1-2\%$ of the users pass the threshold.
 Figure 14: Cross-comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$.